### v1

# Robust $\Lambda$-quantiles and extreme probabilities

Abstract
In this paper, we investigate the robust models for $\Lambda$-quantiles with partial information regarding the loss distribution, where $\Lambda$-quantiles extend the classical quantiles by replacing the fixed probability level with a probability/loss function $\Lambda$. We find that, under some assumptions, the robust $\Lambda$-quantiles equal the $\Lambda$-quantiles of the extreme probabilities. This finding allows us to obtain the robust $\Lambda$-quantiles by applying the results of robust quantiles in the literature. Our results are applied to uncertainty sets characterized by three different constraints respectively: moment constraints, probability distance constraints via Wasserstein metric, and marginal constraints in risk aggregation. We obtain some explicit expressions for robust $\Lambda$-quantiles by deriving the extreme probabilities for each uncertainty set. Those results are applied to optimal portfolio selection under model uncertainty.
Key-words: Quantiles; $\Lambda$-quantiles; Distributionally robust optimization; Robust $\Lambda$-quantiles; Extreme probabilities; Wasserstein distance; Portfolio selection
1 Introduction
To assess the impact of model misspecification and offer a robust quantification of the optimization problem, distributionally robust optimization (DRO) has been studied extensively recently and become a fast-growing field. In DRO, some object of interest is assessed over a set of alternative distributions, where the object can be some risk measures (functionals) or expected utility and the set of alternative distributions are called the uncertainty set or ambiguity set. Typically, we are concerned with the worst-case and best-case values of some risk measures over the uncertainty set, called the robust risk measures, representing the robust quantification of the risk measures, which is independent of the choice of the distributions lying in the uncertainty set. The importance of robustness of risk measures has been emphasized in the academic response of Basel Accord 3.5 in Embrechts et al. (2014), and the robustness of the risk measures has been studied in e.g., Cont et al. (2010) and He et al. (2024).
The uncertainty sets are always characterized by some constraints, representing the partially known information. The moment constraints are popularly used to characterize the uncertainty sets in the DRO literature. It means some information on the moments of the distributions is known. Under this type of uncertainty sets, robust-quantiles were studied in Ghaoui et al. (2003), robust expected shortfall (ES) was considered in Natarajan et al. (2010) and Zhu and Fukushima (2009), robust convex risk measures were investigated in Li (2018), and robust distortion risk measures and riskmetrics were considered in Pesenti et al. (2020), Shao and Zhang (2023), Shao and Zhang (2024), and Cai et al. (2023). One popular probability distance constraint is defined by Wasserstein metric, where the Wasserstein distance represents the tolerance of the discrepancy of the alternative distributions with the baseline distribution; see e.g., Esfahani and Kuhn (2018) and Blanchet and Murthy (2019). The worst-case values of quantiles, spectral risk measures and distortion riskmetrics over this uncertainty set were studied in Bernard et al. (2024), Liu et al. (2022) and Pesenti et al. (2020). The marginal constraint is always imposed on the risk aggregation, where the marginal distributions are known but the dependence structure is completely unknown. For this uncertainty set, the robust-quantiles are the main concern in the literature, which is only solved for some special marginal distributions. The robust-quantiles were derived for marginal distributions with monotone densities in Wang and Wang (2016), Jakobsons et al. (2016) and Blanchet et al. (2020). We refer to Puccetti and Rüschendorf (2012) and Eckstein et al. (2020) for numerical studies. The uncertainty sets characterized by the above three constraints will be explored later.
In this paper, we focus on the $\Lambda$-quantiles, which is an extension of quantiles by replacing the fixed probability level by a probability/loss function. As one of the most important measures of risk in quantitative risk management, quantiles have been criticized by two major reasons: First, it cannot capture the tail risk; Second, it is lack of convexity (or subadditivity), resulting in penalizing diversification. More detailed discussions on it can be found in, e.g., McNeil et al. (2015). To overcome some drawbacks of quantiles, $\Lambda$-quantiles was introduced by Frittelli et al. (2014) as an extension of quantiles. In the recent literature, it shows that $\Lambda$-quantiles possess some desirable properties. From the definition of $\Lambda$-quantiles, it is able to control the tail probability of the risk for increasing $\Lambda$ functions, which was also tested empirically in Hitaj et al. (2018). Moreover, as shown in Han et al. (2024), $\Lambda$-quantiles with decreasing $\Lambda$ functions satisfy quasi-star-shapedness, indicating that convex combination with a determinate risk does not increase the risk, which is a property weaker than quasi-convexity and penalizing the special type of concentration. In addition, $\Lambda$-quantiles with decreasing $\Lambda$ functions also satisfy cash-subadditivity, which is desirable to measure the risk of the financial position in the presence of stochastic or ambiguous interest rates; see El Karoui and Ravanelli (2009) and Han et al. (2024). We refer to Burzoni et al. (2017) for some other properties of $\Lambda$-quantiles such as robustness, elicitability, and consistency.
Typically, the $\Lambda$ function is chosen to be either increasing or decreasing, representing the risk appetite of the decision maker (DM) as shown in Frittelli et al. (2014), and the choice and estimation of the $\Lambda$ function were studied in Hitaj et al. (2018). Recently, an axiomatization of $\Lambda$-quantiles was offered in Bellini and Peri (2022) and a new expression of $\Lambda$-quantiles with decreasing $\Lambda$ function was provided in Han et al. (2024). Additionally, Ince et al. (2022) focused on the risk contribution (sensitivity analysis) of $\Lambda$-quantiles, Balbás et al. (2023) studied the applications of $\Lambda$-quantiles to reinsurance contracts and premium calculation, Liu (2024) investigated the risk sharing problem among multiple agents using $\Lambda$-quantiles to represent their preferences, and Pesenti and Vanduffel (2024) used the scoring function of $\Lambda$-quantiles for elicitability as the cost function to study the Monge-Kantorovich optimal transport.
In this paper, we aim to establish the robust models for $\Lambda$-quantiles with limited information on the underlying distribution of loss, where the function $\Lambda$ can be either decreasing or increasing. Despite the prevailing notion in existing literature that most researchers consider the risk in the worst-case scenario, wherein the DM is extremely ambiguity-averse, empirical studies (see e.g., Heath and Tversky (1991) and Kocher et al. (2015)) demonstrate that the DM’s attitude towards ambiguity is not uniformly negative. In fact, they may exhibit ambiguity-loving tendencies if they perceive themselves as knowledgeable or competent in the context. Thus, both the worst and best possible values for $\Lambda$-quantiles over a set of alternative distributions will be considered.
In Section 2, we introduce four different definitions of $\Lambda$-quantiles given by Bellini and Peri (2022) and summarize some new properties of $\Lambda$-quantiles. In particular, we show the quasi-star-shapedness for all four $\Lambda$-quantiles with decreasing $\Lambda$ functions and show that all those four $\Lambda$-quantiles are cash-subadditive for decreasing $\Lambda$ functions and cash-supadditive for increasing $\Lambda$ functions. Those properties motivate the study and application of $\Lambda$-quantiles with monotone $\Lambda$ functions. Moreover, we offer an alternative expression for one $\Lambda$-quantile with decreasing $\Lambda$ functions as an analog of the expression in Han et al. (2024), which enhances interpretability and offers convenience for computing robust $\Lambda$-quantiles.
In Section 3, we show that, under some assumption, the robust $\Lambda$-quantiles over a general uncertainty set is equal to the $\Lambda$-quantile of the extreme probabilities over the same uncertainty set, which is the main finding of this paper. This is an unexpected result. This fact shows that the computation of robust $\Lambda$-quantiles can be decomposed as two steps: i) compute the extreme probabilities; ii) compute the $\Lambda$-quantiles of the obtained extreme probabilities. It means that in order to compute the robust $\Lambda$-quantiles, it suffices to compute the extreme probabilities, which can be derived using the results of robust quantiles. The crucial assumption for our results is the attainability of extreme probabilities. Verifying this property for popular uncertainty sets is not trivial, and will be discussed in Section 4. Moreover, we construct several counterexamples for the assumptions of our main results and discuss the existence of the optimal distributions in the robust $\Lambda$-quantiles. It is worth mentioning that our findings are closely liked to Mao et al. (2024), where a novel approach, called model aggregation approach, was introduced to evaluate the risk under model uncertainty for some risk measures including quantiles and ES.
In Section 4, we consider the uncertainty sets characterized by three different constraints respectively: moment constraints, probability constraints via Wasserstein distance, and marginal constraints in risk aggregation. We obtain the extreme probabilities for each uncertainty set replying on the results on robust quantiles over the same uncertainty sets in the literature and check the attainability of the extreme probabilities for each uncertainty set. Applying our main results in Section 3, we directly obtain the robust $\Lambda$-quantiles, showing the usefulness of our findings. Moreover, we apply our results to the portfolio selection problem with model uncertainty, where the risk of the portfolio is assessed by $\Lambda$-quantiles. In particular, for the uncertainty sets with marginal constraints in risk aggregation, we show that diversification is penalized for the portfolio of assets with the same marginal distribution, extending the result for quantiles in Chen et al. (2022). Some numerical studies have been done to show the values of the robust $\Lambda$-quantiles and optimal portfolio positions. The paper is concluded in Section 5.
2 Notation and preliminaries
Let $\mathcal{X}$ be a set of random variables in a given atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$. Denote by $\mathcal{M}$ the set of cumulative distribution functions (cdfs) of all random variables in $\mathcal{X}$. For any law-invariant risk measures, it can be defined on either $\mathcal{X}$ or $\mathcal{M}$, which makes no difference for the understanding. For a distribution $F\in\mathcal{M}$, its left quantile at level $\alpha\in[0,1]$ is defined by
$q_{\alpha}^{-}(F)=F^{-1}(\alpha)=\inf\{x\in\mathbb{R}:F(x)\geqslant\alpha\},$
and its right quantile at level $\alpha\in[0,1]$ is given by
$q_{\alpha}^{+}(F)=F_{+}^{-1}(\alpha)=\inf\{x\in\mathbb{R}:F(x)>\alpha\}$
with the convention that $\inf\emptyset=\infty$. Note that by the above definition, we have $q_{0}^{-}(F)=-\infty$ and $q_{1}^{+}(F)=\infty$. We refer to e.g., McNeil et al. (2015) for more discussions on the properties and application of quantiles. We next present four definitions of $\Lambda$-quantiles given in Bellini and Peri (2022).
Definition 1.
For $\Lambda:\mathbb{R}\to[0,1]$, and an increasing function $f:\mathbb{R}\to[0,1]$, the $\Lambda$-quantiles of $f$ are defined as [^1] Note that we use the notation $f$ here to distinguish it from $F$, as $f$ is not necessarily a distribution function.
$q_{\Lambda}^{-}(f)=\inf\{x\in\mathbb{R}:f(x)\geqslant\Lambda(x)\},$ $~{}~{}q_{\Lambda}^{+}(f)=\inf\{x\in\mathbb{R}:f(x)>\Lambda(x)\},$
$\tilde{q}_{\Lambda}^{-}(f)=\sup\{x\in\mathbb{R}:f(x)<\Lambda(x)\},$ $~{}~{}\tilde{q}_{\Lambda}^{+}(f)=\sup\{x\in\mathbb{R}:f(x)\leqslant\Lambda(x)\},$ (1)
where $\inf\emptyset=\infty$ and $\sup\emptyset=-\infty$ by convention.
In Definition 1, agents can choose $\Lambda$ as either increasing or decreasing functions. Throughout this paper, “increasing” (resp. “decreasing”) means non-decreasing (resp. non-increasing). Without additional assumptions, the four $\Lambda$-quantiles can all differ; see Examples 4 and 5 of Bellini and Peri (2022). In general, an increasing $\Lambda$ implies that the DM only accepts smaller probability for larger losses. This reflects a relatively more risk-averse attitude. Conversely, a decreasing $\Lambda$ suggests a tendency to tolerate a larger probability as losses increase. This indicates that the DM is more relatively risk-seeking.
A simple example of $\Lambda$-quantile is the two-level $\Lambda$-quantile presented in Example 7 of Bellini and Peri (2022), where the $\Lambda$ function is defined as $\Lambda:x\mapsto\beta\mathds{1}_{\{x<z\}}+\alpha\mathds{1}_{\{x\geqslant z\}}$. In particular, if $\Lambda$ is a constant, then $\Lambda$-quantiles boil down to the quantiles.
A mapping $\rho:\mathcal{X}\rightarrow\mathbb{R}$ is said to satisfy monotonicity if $\rho(X)\leqslant\rho(Y)$ for all $X,Y\in\mathcal{X}$ with $X\leqslant Y$, and satisfy cash additivity if $\rho(X+m)=\rho(X)+m$ for all $X\in\mathcal{X}$ and $m\in\mathbb{R}$. It is well known that $q^{-}_{\alpha}$ and $q^{+}_{\alpha}$ satisfy the both two properties. However, $\Lambda$-quantiles are not cash-additive in general.
A mapping $\rho:\mathcal{X}\rightarrow\mathbb{R}$ is said to satisfy cash subadditivity (resp. cash supadditivity) if $\rho(X+m)\leqslant\rho(X)+m$ (resp. $\rho(X+m)\geqslant\rho(X)+m$) for all $X\in\mathcal{X}$ and $m\geqslant 0$. A mapping $\rho$ is said to be quasi-star-shapedness if $\rho(\lambda X+(1-\lambda)t)\leqslant\max\{\rho(X),\rho(t)\}$ for all $X\in\mathcal{X},t\in\mathbb{R}$ and $\lambda\in[0,1]$. In the context of capital requirement, cash subadditivity or supadditivity allows a non-linear increase of the capital requirement as cash is added to the future financial position. Cash-subadditivity is desirable if the interest rate is stochastic or ambiguous, and it allows the current reserve and the future risk positions to preserve their own numéraire. Here cash-subadditivity reflects the time value of money; see El Karoui and Ravanelli (2009) and Han et al. (2024). Quasi-star-shapedness means that $\rho$ has quasi-convexity at each constant, representing some consideration of diversification, which is weaker than quasi-convexity. Its theoretic-decision interpretation is available at Han et al. (2024).
In Theorem 3.1 of Han et al. (2024), a representation of $q_{\Lambda}^{-}$ is given in terms of quantiles. In the following proposition, we obtain a similar formula for $q_{\Lambda}^{+}$. For completeness, we also include the representation for $q_{\Lambda}^{-}$ in the following proposition.
Proposition 1.
If $\Lambda:\mathbb{R}\to[0,1]$ is a decreasing function, then for $F\in\mathcal{M}$, we have
$q_{\Lambda}^{-}(F)=\inf_{x\in\mathbb{R}}\left\{q_{\Lambda(x)}^{-}(F)\vee x\right\}=\sup_{x\in\mathbb{R}}\left\{q_{\Lambda(x)}^{-}(F)\wedge x\right\},$ (2)
and
$q_{\Lambda}^{+}(F)=\inf_{x\in\mathbb{R}}\left\{q_{\Lambda(x)}^{+}(F)\vee x\right\}=\sup_{x\in\mathbb{R}}\left\{q_{\Lambda(x)}^{+}(F)\wedge x\right\}.$ (3)
Proof.
Note that (2) follows directly from Theorem 3.1 of Han et al. (2024). Next, we show (3). Note that for $x\in\mathbb{R}$ and $t\in[0,1]$, $F(x)\leqslant t$ implies $q^{+}_{t}(F)\geqslant x$. Hence, for a decreasing $\Lambda$, we have
$q_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(F)$ $=\sup\{x\in\mathbb{R}:F(x)\leqslant\Lambda(x)\}$
$\leqslant\sup\{x\in\mathbb{R}:q_{\Lambda(x)}^{+}(F)\geqslant x\}$
$=\sup\{q_{\Lambda(x)}^{+}(F)\wedge x:q_{\Lambda(x)}^{+}(F)\geqslant x\}\leqslant\sup_{x\in\mathbb{R}}\left\{q_{\Lambda(x)}^{+}(F)\wedge x\right\}.$
Moreover, using the fact $F(x)>t$ implies $q^{+}_{t}(F)\leqslant x$, we have
$q_{\Lambda}^{+}(F)$ $=\inf\{x\in\mathbb{R}:F(x)>\Lambda(x)\}$
$\geqslant\inf\{x\in\mathbb{R}:q_{\Lambda(x)}^{+}(F)\leqslant x\}$
$=\inf\{q_{\Lambda(x)}^{+}(F)\vee x:q_{\Lambda(x)}^{+}(F)\leqslant x\}\geqslant\inf_{x\in\mathbb{R}}\left\{q_{\Lambda(x)}^{+}(F)\vee x\right\}.$
Since $q_{\Lambda(x)}^{+}(F)\wedge x\leqslant q_{\Lambda(y)}^{+}(F)\vee y$ for any $x,y\in\mathbb{R}$, we have
$q_{\Lambda}^{+}(F)\leqslant\sup_{x\in\mathbb{R}}\left\{q_{\Lambda(x)}^{+}(F)\wedge x\right\}\leqslant\inf_{x\in\mathbb{R}}\left\{q_{\Lambda(x)}^{+}(F)\vee x\right\}\leqslant q_{\Lambda}^{+}(F),$
which implies (3). ∎
In Bellini and Peri (2022), many properties of $\Lambda$-quantiles are discussed such as monotonicity, locality, and quasi-convexity on distributions. We next summarize some of the properties possessed by $\Lambda$-quantiles that were not included in Bellini and Peri (2022) in the following proposition.
Proposition 2.
The following statements hold.
  • (i)
    $q_{\Lambda}^{-}(F)\leqslant\tilde{q}_{\Lambda}^{-}(F)$ and $q_{\Lambda}^{+}(F)\leqslant\tilde{q}_{\Lambda}^{+}(F)$, and the inequalities become equalities if $\Lambda$ is decreasing;
  • (ii)
    All the four $\Lambda$-quantiles $q_{\Lambda}^{-},q_{\Lambda}^{+},\tilde{q}_{\Lambda}^{-}$, and $\tilde{q}_{\Lambda}^{+}$ are cash-subadditive if $\Lambda$ is decreasing and cash-supadditive if $\Lambda$ is increasing;
  • (iii)
    If $\Lambda$ is decreasing, all the four $\Lambda$-quantiles $q_{\Lambda}^{-},q_{\Lambda}^{+},\tilde{q}_{\Lambda}^{-}$, and $\tilde{q}_{\Lambda}^{+}$ are quasi-star-shaped.
Proof.
We first focus on (i). We first show $q_{\Lambda}^{-}(F)\leqslant\tilde{q}_{\Lambda}^{-}(F)$. Clearly, if $\tilde{q}_{\Lambda}^{-}(F)=-\infty$, then $F(x)\geqslant\Lambda(x)$ for all $x\in\mathbb{R}$. This implies $q_{\Lambda}^{-}(F)=-\infty$. Hence, $q_{\Lambda}^{-}(F)\leqslant\tilde{q}_{\Lambda}^{-}(F)$ holds. If $\tilde{q}_{\Lambda}^{-}(F)=+\infty$, the inequality holds trivially. Next, we consider the case $x_{0}=\tilde{q}_{\Lambda}^{-}(F)\in\mathbb{R}$. By the definition, we have $F(x)\geqslant\Lambda(x)$ for all $x>x_{0}$, which implies $q_{\Lambda}^{-}(F)\leqslant x_{0}=\tilde{q}_{\Lambda}^{-}(F)$. We can similarly show $q_{\Lambda}^{+}(F)\leqslant\tilde{q}_{\Lambda}^{+}(F)$. The equalities for decreasing $\Lambda$ follows directly from Proposition 6 of Bellini and Peri (2022).
Next, we show (ii). Note that for $c\geqslant 0$, $q_{\Lambda}^{-}(X+c)=q^{-}_{\Lambda^{c}}(X)+c$, where $\Lambda^{c}(t)=\Lambda(t+c)$ for $t\in\mathbb{R}$. Moreover, $\Lambda^{c}\leqslant\Lambda$ if $\Lambda$ is decreasing and $\Lambda^{c}\geqslant\Lambda$ if $\Lambda$ is increasing. By Proposition 3 of Bellini and Peri (2022), we have $q_{\Lambda^{c}}^{-}(X)\leqslant q_{\Lambda}^{-}(X)$ if $\Lambda$ is decreasing, and $q_{\Lambda^{c}}^{-}(X)\geqslant q_{\Lambda}^{-}(X)$ if $\Lambda$ is increasing. Therefore, if $\Lambda$ is decreasing, $q_{\Lambda}^{-}(X+c)\leqslant q_{\Lambda}^{-}(X)+c$, implying $q_{\Lambda}^{-}$ is cash-subadditive; if $\Lambda$ is increasing, $q_{\Lambda}^{-}(X+c)\geqslant q_{\Lambda}^{-}(X)+c$, implying $q_{\Lambda}^{-}$ is cash-supadditive. The analogous arguments can be employed to show the properties for other $\Lambda$-quantiles.
Finally, we consider (iii). By (i), we have $\tilde{q}_{\Lambda}^{-}=q_{\Lambda}^{-}$. Hence, it follows from Theorem 3.1 of Han et al. (2024) that $\tilde{q}_{\Lambda}^{-}$ and $q_{\Lambda}^{-}$ are quasi-star-shaped. Similarly, in light of (i), we have $\tilde{q}_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(F)$. Thus it suffices to show that $q_{\Lambda}^{+}$ is quasi-star-shaped. We first show $q_{\alpha}^{+}\vee x$ is quasi-star-shaped for $\alpha\in[0,1]$ and $x\in\mathbb{R}$. If $\alpha=1$, it is trivial. Then we suppose $\alpha\in[0,1)$. For $\lambda\in[0,1],t\in\mathbb{R}$ and $X\in\mathcal{X}$, we have
$q_{\alpha}^{+}(\lambda X+(1-\lambda)t)\vee x=\left(\lambda q_{\alpha}^{+}(X)+(1-\lambda)t\right)\vee x\leqslant\max(q_{\alpha}^{+}(X)\vee x,t\vee x).$
Hence, $q_{\alpha}^{+}\vee x$ is quasi-star-shaped. In light of Lemma 3.1 of Han et al. (2024) and the expression (3), we have $q_{\Lambda}^{+}$ is quasi-star-shaped. ∎
Cash supadditivity is the dual property of cash subadditivity. As we can see from (ii) of Proposition 2, whether a $\Lambda$-quantile satisfies cash subadditivity or cash supadditivity depends on the direction of monotonicity of the $\Lambda$ function. The intermediate property is cash additivity. A $\Lambda$-quantile with monotone $\Lambda$ functions satisfies cash additivity if and only if it is a quantile; see Proposition 1 of Liu (2024).
Note that $q_{\Lambda}^{-}$ or $q_{\Lambda}^{+}$ with increasing $\Lambda$ is in general not quasi-star-shaped even when $\Lambda$ is a continuous function. One can see the following counterexample.
Example 1.
For $10<\alpha<1/2<\beta<$ 1, consider the following continuous and increasing function
$\Lambda(x)=\alpha\mathds{1}_{\{x\leqslant 1/2\}}+((\beta-\alpha)x-\beta/2+3\alpha/2)\mathds{1}_{\{1/2<x<3/2\}}+\beta\mathds{1}_{\{x\geqslant 3/2\}},~{}~{}x\in\mathbb{R}.$
For $t=7/4$, $\lambda=1/8$, and a random variable $X$ satisfying $\mathbb{P}(X=0)=\mathbb{P}(X=2)=1/2$, we have
$\mathbb{P}(\lambda X+(1-\lambda)t=57/32)=\mathbb{P}(\lambda X+(1-\lambda)t=49/32)=1/2.$
Direct computation gives
$q_{\Lambda}^{-}(X)=0,~{}q_{\Lambda}^{-}(\lambda X+(1-\lambda)t)=57/32,~{}q_{\Lambda}^{-}(t)=7/4,$
which imply $q_{\Lambda}^{-}(\lambda X+(1-\lambda)t)>\max\{q_{\Lambda}^{-}(X),q_{\Lambda}^{-}(t)\}.$ Hence, $q_{\Lambda}^{-}$ is not quasi-star-shaped. Note that all the above arguments still hold if we replace $q_{\Lambda}^{-}$ by $q_{\Lambda}^{+}$. Hence, $q_{\Lambda}^{+}$ is not quasi-star-shaped.
3 Robust $\Lambda$-quantiles
In this section, we investigate the worst-case and best-case values of $\Lambda$-quantiles over general uncertainty sets, which is our primary objective in this paper. For a set of distributions $\mathcal{M}$, we denote $F_{\mathcal{M}}^{-}(x)=\inf_{F\in\mathcal{M}}F(x)$, $x\in\mathbb{R}$ and $F_{\mathcal{M}}^{+}(x)=\sup_{F\in\mathcal{M}}F(x)$, $x\in\mathbb{R}$. Note that both $F_{\mathcal{M}}^{-}$ and $F_{\mathcal{M}}^{+}$ are increasing functions. Additionally, $F_{\mathcal{M}}^{-}$ is right-continuous, whereas $F_{\mathcal{M}}^{+}$ may be not. Both of them may not be distribution functions. We say that $F_{\mathcal{M}}^{-}$ (resp. $F_{\mathcal{M}}^{+}$) is attainable if for any $z\in\mathbb{R}$, there exists $F\in\mathcal{M}$ such that $F(z)=F_{\mathcal{M}}^{-}(z)$ (resp. $F(z)=F_{\mathcal{M}}^{+}(z)$). The following theorem is the main finding of this paper.
Theorem 1.
Let $\mathcal{M}$ be a set of distributions and $\Lambda:\mathbb{R}\to[0,1]$. Then we have
  1. (i)
    $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})$ and $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(F_{\mathcal{M}}^{+})$;
  2. (ii)
    $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})$ if $F_{\mathcal{M}}^{-}$ is attainable, and $\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(F_{\mathcal{M}}^{+})$ if $F_{\mathcal{M}}^{+}$ is attainable;
  3. (iii)
    if $\Lambda$ is decreasing, then (i) remains true by replacing $\tilde{q}_{\Lambda}^{-}$ by $q_{\Lambda}^{-}$ and $q_{\Lambda}^{+}$ by $\tilde{q}_{\Lambda}^{+}$ and (ii) remains true by replacing $\tilde{q}_{\Lambda}^{+}$ by $q_{\Lambda}^{+}$ and $q_{\Lambda}^{-}$ by $\tilde{q}_{\Lambda}^{-}$.
Proof.
Case (i). We start with the first equality. Note that $F_{\mathcal{M}}^{-}(x)\leqslant F(x),~{}x\in\mathbb{R}$ for any $F\in\mathcal{M}$. Hence in light of the monotonicity of $\Lambda$-quantile, we have $\tilde{q}_{\Lambda}^{-}(F)\leqslant\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})$ for all $F\in\mathcal{M}$. This implies $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)\leqslant\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})$.
We next show the inverse inequality. Let $z=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})$. If $z=-\infty$, $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})=-\infty$ holds obviously. If $z=\infty$, then there exists a sequence of $x_{n}$ such that $x_{n}\uparrow\infty$ and $F_{\mathcal{M}}^{-}(x_{n})<\Lambda(x_{n})$. For any fixed $x_{n}$, we could find $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{-}(x_{n})\leqslant G(x_{n})<\Lambda(x_{n})$. This implies $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)\geqslant\tilde{q}_{\Lambda}^{-}(G)\geqslant x_{n}\to\infty$ as $n\to\infty$. Hence, it follows that $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})$.
Next, we suppose $z\in\mathbb{R}$. If $F_{\mathcal{M}}^{-}(z)<\Lambda(z)$, there exists $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{-}(z)\leqslant G(z)<\Lambda(z)$. Hence, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)\geqslant\tilde{q}_{\Lambda}^{-}(G)\geqslant z$. If $F_{\mathcal{M}}^{-}(z)\geqslant\Lambda(z)$, there exists a sequence $y_{n}\uparrow z$ such that $F_{\mathcal{M}}^{-}(y_{n})<\Lambda(y_{n})$. For any fixed $y_{n}$, there exists a distribution $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{-}(y_{n})\leqslant G(y_{n})<\Lambda(y_{n})$. This means $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)\geqslant\tilde{q}_{\Lambda}^{-}(G)\geqslant y_{n}$. Letting $n\to\infty$, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)\geqslant z$. We establish the inverse inequality.
We now focus on the second equality. Observe that $F_{\mathcal{M}}^{+}(x)\geqslant F(x),~{}x\in\mathbb{R}$ for any $F\in\mathcal{M}$. It follows from the monotonicity of $\Lambda$-quantile that $q_{\Lambda}^{+}(F_{\mathcal{M}}^{+})\leqslant q_{\Lambda}^{+}(F)$ for all $F\in\mathcal{M}$. Consequently, $q_{\Lambda}^{+}(F_{\mathcal{M}}^{+})\leqslant\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)$.
Let $z=q_{\Lambda}(F_{\mathcal{M}}^{+})$. If $z=\infty$, then $q_{\Lambda}^{+}(F_{\mathcal{M}}^{+})=\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)$ holds trivially. If $z=-\infty$, there exists a sequence of $x_{n}\downarrow-\infty$ such that $F_{\mathcal{M}}^{+}(x_{n})>\Lambda(x_{n})$. For any fixed $x_{n}$, there exists $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{+}(x_{n})\geqslant G(x_{n})>\Lambda(x_{n})$. This implies $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)\leqslant q_{\Lambda}^{+}(G)\leqslant x_{n}\to-\infty$ as $n\to\infty$. Hence, we obtain $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)=q_{\Lambda}(F_{\mathcal{M}}^{+})$.
We suppose $z\in\mathbb{R}$. If $F_{\mathcal{M}}^{+}(z)>\Lambda(z)$, there exists $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{+}(z)\geqslant G(z)>\Lambda(z)$, implying $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)\leqslant q_{\Lambda}^{+}(G)\leqslant z$. If $F_{\mathcal{M}}^{+}(z)\leqslant\Lambda(z)$, there exists a sequence of $y_{n}\downarrow z$ such that $F_{\mathcal{M}}^{+}(y_{n})>\Lambda(y_{n})$. For fixed $y_{n}$, there exists $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{+}(y_{n})\geqslant G(y_{n})>\Lambda(y_{n})$, which implies $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)\leqslant q_{\Lambda}^{+}(G)\leqslant y_{n}\to z$ as $n\to\infty$. We establish the inverse inequality.
Case (ii). For the first equality, note that $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)\leqslant\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})$ follows from the same reasoning as in the proof of case (i). We only need to show the inverse inequality.
Denote $\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})$ by $z$. If $z=-\infty$, then $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})=-\infty$ holds obviously. If $z=\infty$, then there exists a sequence of $x_{n}$ such that $x_{n}\uparrow\infty$ and $F_{\mathcal{M}}^{-}(x_{n})\leqslant\Lambda(x_{n})$. For any fixed $x_{n}$, using the attainability of $F_{\mathcal{M}}^{-}$, there exists $G\in\mathcal{M}$ such that $G(x_{n})=F_{\mathcal{M}}^{-}(x_{n})\leqslant\Lambda(x_{n})$. This implies $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)\geqslant\tilde{q}_{\Lambda}^{+}(G)\geqslant x_{n}\to\infty$ as $n\to\infty$. Hence, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})$.
Next, we suppose $z\in\mathbb{R}$. If $F_{\mathcal{M}}^{-}(z)\leqslant\Lambda(z)$, there exists $G\in\mathcal{M}$ such that $G(z)=F_{\mathcal{M}}^{-}(z)\leqslant\Lambda(z)$. Hence, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)\geqslant\tilde{q}_{\Lambda}^{+}(G)\geqslant z$. If $F_{\mathcal{M}}^{-}(z)>\Lambda(z)$, there exists a sequence $y_{n}\uparrow z$ such that $F_{\mathcal{M}}^{-}(y_{n})\leqslant\Lambda(y_{n})$. For any fixed $y_{n}$, there exists a distribution $G\in\mathcal{M}$ such that $G(y_{n})=F_{\mathcal{M}}^{-}(y_{n})\leqslant\Lambda(y_{n})$. This means $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)\geqslant\tilde{q}_{\Lambda}^{+}(G)\geqslant y_{n}$. Letting $n\to\infty$, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)\geqslant z$. We establish the inverse inequality.
The proof of the second equality of (ii) follows the similar argument as that of the second equality of Case (i) by applying the attainability of $F_{\mathcal{M}}^{-}$. We omit the details.
Case (iii). If $\Lambda$ is decreasing, by (i) of Proposition 2, we have $q_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F)$ and $q_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(F)$. The conclusion in (iii) is directly implied by the conclusions in (i)-(ii). ∎
Theorem 1 shows that finding the robust $\Lambda$-quantiles can be decomposed into two steps: First, find the supremum or infimum value of probabilities, i.e., $F_{\mathcal{M}}^{+}$ and $F_{\mathcal{M}}^{-}$; Second, compute the $\Lambda$-quantiles of those extreme probabilities. The second step is indeed trivial and the first step is the main step, which is a problem in probability theory with a long history; see Makarov (1981) and Rüschendorf (1982) for early results, and see Li (2018), Blanchet et al. (2020) and Pesenti et al. (2020) for the recent results. Surprisingly, the complexity of the $\Lambda$ functions does not significantly increase any difficulty to compute the robust $\Lambda$-quantiles. This is unexpected! Roughly speaking, the main message of our results is that if we know the extreme probabilities, applying Theorem 1, we could obtain the robust $\Lambda$-quantiles immediately. This enables us to obtain the robust $\Lambda$-quantiles with the aid of many results on robust-quantiles in the literature.
Applying Theorem 1, we obtain the following conclusion for robust quantiles, which indicates the relation between robust-quantiles and extreme probabilities.
Corollary 1.
For $\alpha\in(0,1)$, we have
  1. (i)
    $\sup_{F\in\mathcal{M}}q_{\alpha}^{-}(F)=q_{\alpha}^{-}(F_{\mathcal{M}}^{-})$ and $\inf_{F\in\mathcal{M}}q_{\alpha}^{+}(F)=q_{\alpha}^{+}(F_{\mathcal{M}}^{+})$;
  2. (ii)
    $\sup_{F\in\mathcal{M}}q_{\alpha}^{+}(F)=q_{\alpha}^{+}(F_{\mathcal{M}}^{-})$ if $F_{\mathcal{M}}^{-}$ is attainable, and $\inf_{F\in\mathcal{M}}q_{\alpha}^{-}(F)=q_{\alpha}^{-}(F_{\mathcal{M}}^{+})$ if $F_{\mathcal{M}}^{+}$ is attainable.
Note that the conclusion in (iii) of Theorem 1 requires an additional assumption that $\Lambda$ is decreasing. In general, the conclusion in (iii) of Theorem 1 does not hold for other $\Lambda$ functions even for increasing $\Lambda$ functions. We can see this from the following example.
Example 2.
Let
$\Lambda(x)=\left\{\begin{array}[]{cc}1/4,&x<0\\ (1+x)/4,&0\leqslant x<1\\ 1/2,&x\geqslant 1\end{array}\right.,$
and =$\mathcal{M}=\{F_{1},F_{2}\}$ with
$F_{1}(x)=\left\{\begin{array}[]{cc}0,&x<0\\ 1/3,&0\leqslant x<1\\ 1,&x\geqslant 1\end{array}\right.,~{}\text{and}~{}F_{2}(x)=\left\{\begin{array}[]{cc}0,&x<1/2\\ 1,&x\geqslant 1/2\end{array}\right..$
Direct computation gives
$F_{\mathcal{M}}^{-}(x)=\left\{\begin{array}[]{cc}0,&x<1/2\\ 1/3,&1/2\leqslant x<1\\ 1,&x\geqslant 1\end{array}\right.,$
$q_{\Lambda}^{-}(F_{1})=0$, $q_{\Lambda}^{-}(F_{2})=1/2$ and $q_{\Lambda}^{-}(F_{\mathcal{M}}^{-})=1$. Hence $\sup_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)=1/2<1=q_{\Lambda}^{-}(F_{\mathcal{M}}^{-})$. We can similarly construct counterexamples for other three $\Lambda$-quantiles in (iii) of Theorem 1 for increasing $\Lambda$ functions.
Note also that it requires the attainability of $F_{\mathcal{M}}^{-}$ and $F_{\mathcal{M}}^{+}$ in (ii) of Theorem 1. If we have finite candidates in $\mathcal{M}$, i.e., =$\mathcal{M}=\{F_{1},\dots,F_{n}\}$. Then $F_{\mathcal{M}}^{-}(x)=\bigwedge_{i=1}^{n}F_{i}(x)$ and $F_{\mathcal{M}}^{+}(x)=\bigvee_{i=1}^{n}F_{i}(x)$, and both $F_{\mathcal{M}}^{-}$ and $F_{\mathcal{M}}^{+}$ are attainable, where $\bigwedge_{i=1}^{n}x_{i}=\min(x_{1},\dots,x_{n})$ and $\bigvee_{i=1}^{n}x_{i}=\max(x_{1},\dots,x_{n})$. Applying Theorem 1, we arrive at the following results.
Corollary 2.
Let =$\mathcal{M}=\{F_{1},\dots,F_{n}\}$ and $\Lambda:\mathbb{R}\to[0,1]$. Then we have
  1. (i)
    $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(\bigwedge_{i=1}^{n}F_{i})$, $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(\bigvee_{i=1}^{n}F_{i})$, $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(\bigwedge_{i=1}^{n}F_{i})$, and $\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(\bigvee_{i=1}^{n}F_{i})$;
  2. (ii)
    if $\Lambda$ is decreasing, then $\sup_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(\bigwedge_{i=1}^{n}F_{i})$, $\inf_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(\bigvee_{i=1}^{n}F_{i})$, $\sup_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(\bigwedge_{i=1}^{n}F_{i})$, and $\inf_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(\bigvee_{i=1}^{n}F_{i})$.
The property of attainability is crucial in the proof of (ii) of Theorem 1 and it is not redundant if the candidates in $\mathcal{M}$ is infinite. We have the following counterexample.
Example 3.
Let
$\Lambda(x)=\left\{\begin{array}[]{cc}3/4,&x<0\\ 1/2,&0\leqslant x<1\\ 1/4,&x\geqslant 1\end{array}\right.,$
and =$\mathcal{M}=\{F_{n},n\geqslant 1\}$ with
$F_{n}(x)=\left\{\begin{array}[]{cc}0,&x<0\\ 1/2+1/(2n),&0\leqslant x<1\\ 1,&x\geqslant 1\end{array}\right..$
Direct computation gives
$F_{\mathcal{M}}^{-}(x)=\left\{\begin{array}[]{cc}0,&x<0\\ 1/2,&0\leqslant x<1\\ 1,&x\geqslant 1\end{array}\right.,$
$\tilde{q}_{\Lambda}^{+}(F_{n})=0,~{}n\geqslant 1$ and $\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})=1$. Hence, $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)<\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})$. We can similarly construct examples to show that $\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)>q_{\Lambda}^{-}(F_{\mathcal{M}}^{+})$ if $F_{\mathcal{M}}^{+}$ is not attainable.
The above example shows that if the attainability fails to hold, then (ii) of Theorem 1 may not be valid. However, we still have the following conclusion:
$\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)\leqslant\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})~{}\text{and}~{}\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)\geqslant q_{\Lambda}^{-}(F_{\mathcal{M}}^{+}).$
Hence, both $\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})$ and $q_{\Lambda}^{-}(F_{\mathcal{M}}^{+})$ are still useful by offering the upper and lower bounds of the corresponding robust $\Lambda$-quantiles.
Compared to the attainability of $F_{\mathcal{M}}^{+}$, the attainability of $F_{\mathcal{M}}^{-}$ may not be easy to check as it is difficult to employ the limit argument; see Section 4. Next, we provide an alternative assumption that is much easier to check if $\Lambda$ is decreasing. Let $\widehat{F}_{\mathcal{M}}^{-}(x)=\inf_{F\in\mathcal{M}}F(x-)$ for all $x\in\mathbb{R}$. We say $\widehat{F}_{\mathcal{M}}^{-}$ is attainable if for any $x\in\mathbb{R}$, there exists $F\in\mathcal{M}$ such that $F(x-)=\widehat{F}_{\mathcal{M}}^{-}(x)$.
Proposition 3.
Let $\mathcal{M}$ be a set of distributions and $\Lambda:\mathbb{R}\to[0,1]$. If $\Lambda$ is decreasing and $\widehat{F}_{\mathcal{M}}^{-}$ is continuous and attainable, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})$.
Proof.
Let $\widehat{F}(x)=F(x-)$ for all $x\in\mathbb{R}$. Replacing $F$ by $\widehat{F}$ and $F_{\mathcal{M}}^{-}$ by $\widehat{F}_{\mathcal{M}}^{-}$ in the proof of (ii) of Theorem 1 and using the attainability of $\widehat{F}_{\mathcal{M}}^{-}$, we obtain the conclusion that $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(\widehat{F})=\tilde{q}_{\Lambda}^{+}(\widehat{F}_{\mathcal{M}}^{-})$. Note that for any $\varepsilon>0$, $\widehat{F}_{\mathcal{M}}^{-}(x)\leqslant F_{\mathcal{M}}^{-}(x)\leqslant\widehat{F}_{\mathcal{M}}^{-}(x+\varepsilon)$. Hence, by the continuity of $\widehat{F}_{\mathcal{M}}^{-}$, we have $\widehat{F}_{\mathcal{M}}^{-}=F_{\mathcal{M}}^{-}$. Moreover, as $\Lambda$ is decreasing, we have $\tilde{q}_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(\widehat{F})$. Combing all the above conclusions, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})$. ∎
Based on Proposition 3, we next discuss the existence of the worst or best distribution for the robust $\Lambda$-quantiles.
Proposition 4.
Let $\mathcal{M}$ be a set of distributions and $\Lambda:\mathbb{R}\to[0,1]$. Then we have the following conclusion.
  1. (i)
    If $\Lambda$ is decreasing and left-continuous, and $\widehat{F}_{\mathcal{M}}^{-}$ is continuous and attainable, and $x=\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})\in\mathbb{R}$, then there exists $F^{*}\in\mathcal{M}$ such that $\tilde{q}_{\Lambda}^{+}(F^{*})=\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)$;
  2. (ii)
    If $\Lambda$ is right-continuous, $F_{\mathcal{M}}^{+}$ is attainable and right-continuous and $x=q_{\Lambda}^{-}(F_{\mathcal{M}}^{+})\in\mathbb{R}$, then there exists $F^{*}\in\mathcal{M}$ such that $q_{\Lambda}^{-}(F^{*})=\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)$.
Proof.
(i) In light of Proposition 3, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})=x\in\mathbb{R}$. By the definition, we have either $F_{\mathcal{M}}^{-}(x)\leqslant\Lambda(x)$ or there exists a sequence $x_{n}\uparrow x$ such that $F_{\mathcal{M}}^{-}(x_{n})\leqslant\Lambda(x_{n})$. Note that $F_{\mathcal{M}}^{-}=\widehat{F}_{\mathcal{M}}^{-}$. Hence, $F_{\mathcal{M}}^{-}$ is continuous. For the second case, using the continuity of $F_{\mathcal{M}}^{-}$ and the left-continuity of $\Lambda$, letting $n\to\infty$, we have $F_{\mathcal{M}}^{-}(x)\leqslant\Lambda(x)$. As $\widehat{F}_{\mathcal{M}}^{-}$ is attainable, there exists $F^{*}\in\mathcal{M}$ such that $\widehat{F}^{*}(x)=\widehat{F}_{\mathcal{M}}^{-}(x)=F_{\mathcal{M}}^{-}(x)$ with $\widehat{F}^{*}(x)=F^{*}(x-)$. This implies $\widehat{F}^{*}(x)\leqslant\Lambda(x)$. Hence, $\tilde{q}_{\Lambda}^{+}(\widehat{F}^{*})\geqslant x=\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)$. Note that the decrease of $\Lambda$ implies $\tilde{q}_{\Lambda}^{+}(F^{*})=\tilde{q}_{\Lambda}^{+}(\widehat{F}^{*})$. Hence, we have $\tilde{q}_{\Lambda}^{+}(F^{*})\geqslant\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)$, which implies $\tilde{q}_{\Lambda}^{+}(F^{*})=\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)$.
(ii) First, it follows from Theorem 1 that $\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(F_{\mathcal{M}}^{+})=x\in\mathbb{R}$. By definition, we have either $F_{\mathcal{M}}^{+}(x)\geqslant\Lambda(x)$ or there exists a sequence of $x_{n}\downarrow x$ such that $F_{\mathcal{M}}^{+}(x_{n})\geqslant\Lambda(x_{n})$. For the second case, it follows from the right-continuity of $F_{\mathcal{M}}^{+}$ and $\Lambda$ that $F_{\mathcal{M}}^{+}(x)\geqslant\Lambda(x)$. The attainability of $F_{\mathcal{M}}^{+}$ implies there exists $F^{*}\in\mathcal{M}$ such that $F^{*}(x)=F_{\mathcal{M}}^{+}(x)\geqslant\Lambda(x)$. This implies $q_{\Lambda}^{-}(F^{*})\leqslant x=\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)$. Hence, we have $q_{\Lambda}^{-}(F^{*})=\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)$. ∎
Note that for other $\Lambda$-quantiles, the best or worst case of distribution may not exist, even for the case when $\Lambda=\alpha$. For instance, for the uncertainty set characterized by moment constraints, the optimal distributions for $\sup_{F\in\mathcal{M}}q_{\alpha}^{-}(F)$ and $\inf_{F\in\mathcal{M}}q_{\alpha}^{+}(F)$ do not exist; see e.g., Corollary 4.1 of Bernard et al. (2024).
Finally, we offer alternative expressions of robust $\Lambda$-quantiles using (2) and (3) when $\Lambda$ is decreasing. The expressions can sometimes facilitate the calculation of robust $\Lambda$-quantiles more conveniently; see Sections 4.1 and 4.2 for details.
Proposition 5.
If $\Lambda$ is decreasing, then we have the following representations.
  • (i)
    $\sup_{F\in\mathcal{M}}{q}_{\Lambda}^{-}(F)=\sup_{x\in\mathbb{R}}\{q_{\Lambda(x)}^{-}(F^{-}_{\mathcal{M}})\wedge x\},$ and $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)=\inf_{x\in\mathbb{R}}\{q_{\Lambda(x)}^{+}(F^{+}_{\mathcal{M}})\vee x\}$;
  • (ii)
    $\sup_{F\in\mathcal{M}}{q}_{\Lambda}^{+}(F)=\sup_{x\in\mathbb{R}}\{q_{\Lambda(x)}^{+}(F^{-}_{\mathcal{M}})\wedge x\}$ if $F_{\mathcal{M}}^{-}$ is attainable, and $\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)=\inf_{x\in\mathbb{R}}\{q_{\Lambda(x)}^{-}(F^{+}_{\mathcal{M}})\vee x\}$ if $F_{\mathcal{M}}^{+}$ is attainable.
Proof.
The results directly follow from Theorem 1, and Equations (2)-(3). However, we will present an alternative proof to demonstrate the connection between robust $\Lambda$-quantiles and quantiles. We take the first one as an example. If $\Lambda$ is decreasing, by (2), we have
$\sup_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)$ $=\sup_{F\in\mathcal{M}}\left\{\sup_{x\in\mathbb{R}}\left\{q_{\Lambda(x)}^{-}(F)\wedge x\right\}\right\}$
$=\sup_{x\in\mathbb{R}}\left\{\sup_{F\in\mathcal{M}}\left\{q_{\Lambda(x)}^{-}(F)\wedge x\right\}\right\}=\sup_{x\in\mathbb{R}}\left\{\left\{\sup_{F\in\mathcal{M}}q_{\Lambda(x)}^{-}(F)\right\}\wedge x\right\}.$
It follows from Corollary 1 that $\sup_{F\in\mathcal{M}}q_{\Lambda(x)}^{-}(F)=q_{\Lambda(x)}^{-}(F^{-}_{\mathcal{M}})$ for any $x\in\mathbb{R}$. Hence, we have $\sup_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)=\sup_{x\in\mathbb{R}}\left\{q_{\Lambda(x)}^{-}(F^{-}_{\mathcal{M}})\wedge x\right\}.$
4 Specific uncertainty sets
In this section, our attention is directed towards the uncertainty sets characterized by three distinct and widely used constraints: (a) moment constraints, (b) probability constraints via Wasserstein distance; and (c) marginal constraints in risk aggregation. Furthermore, our results are applied to the portfolio selection problem. Utilizing Theorem 1, we can clearly see that the robust $\Lambda$-quantiles can be derived based on the expressions of robust quantiles. Denote by $\mathcal{M}_{p}$ the set of all cdfs on $\mathbb{R}$ with finite $p$-th moment and $\mathcal{M}_{p}\left(\mathbb{R}^{n}\right)$ the set of all cdfs on $\mathbb{R}^{n}$ with finite $p$-th moment in each component.
4.1 Moment constraints
We first consider the case with only information on the moments of $F\in\mathcal{M}$. For $p>1,m\in\mathbb{R}$, and $v>0$, let
$\mathcal{M}_{p}(m,v)=\left\{F\in\mathcal{M}_{p}:\int_{0}^{1}F^{-1}(t)\mathrm{d}t=m,\int_{0}^{1}|F^{-1}(t)-m|^{p}\mathrm{d}t\leqslant v^{p}\right\}.$ (4)
Note that $\mathcal{M}_{p}(m,v)$ imposes the constraints on the mean and the $p$-th central moment. The worst-case and best-case values of some risk measures over $\mathcal{M}_{p}(m,v)$, especially for $p=2$, have been extensively studied in the literature; see e.g., Ghaoui et al. (2003), Zhu and Fukushima (2009), Chen et al. (2011), Li (2018), Pesenti et al. (2020), Shao and Zhang (2023), Shao and Zhang (2024), Cai et al. (2023) and the references therein.
For any increasing function $g:(0,1)\to\mathbb{R}$, define $g_{+}^{-1}(x)=\inf\{t\in(0,1):g(t)>x\}$ with the convention that $\inf\emptyset=1$. In the following results, we derive the extreme probabilities over the moment uncertainty sets, i.e., $F_{\mathcal{M}_{p}(m,v)}^{-}$ and $F_{\mathcal{M}_{p}(m,v)}^{+}$.
Proposition 6.
For $p>1,m\in\mathbb{R}$ and $v>0$, we have
$F_{\mathcal{M}_{p}(m,v)}^{-}(x)=l_{+}^{-1}(x),~{}~{}F_{\mathcal{M}_{p}(m,v)}^{+}(x)=u_{+}^{-1}(x),~{}x\in\mathbb{R},$ (5)
where, for $\alpha\in(0,1)$,
$l(\alpha)=m+v\alpha\left(\alpha^{p}(1-\alpha)+(1-\alpha)^{p}\alpha\right)^{-1/p},~{}u(\alpha)=m-v(1-\alpha)\left(\alpha^{p}(1-\alpha)+(1-\alpha)^{p}\alpha\right)^{-1/p}.$
In particular, for $p=2$, we have
${F^{-}_{\mathcal{M}_{2}(m,v)}(x)=\frac{(x-m)^{2}}{v^{2}+(x-m)^{2}}\mathds{1}_{\{x\geqslant m\}}},~{}~{}\text{and}~{}~{}F^{+}_{\mathcal{M}_{2}(m,v)}(x)=\frac{v^{2}}{v^{2}+(m-x)^{2}}\mathds{1}_{\{x\leqslant m\}}+\mathds{1}_{\{x>m\}}.$
Moreover, both $\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}$ and $F_{\mathcal{M}_{p}(m,v)}^{+}$ are continuous and attainable.
Proof.
By Corollary 1 of Pesenti et al. (2020), for $\alpha\in(0,1),p>1,m\in\mathbb{R}$ and $v>0$, we have
$\sup_{F\in\mathcal{M}_{p}(m,v)}q_{\alpha}^{-}(F)=m+v\alpha\left(\alpha^{p}(1-\alpha)+(1-\alpha)^{p}\alpha\right)^{-1/p}:=l(\alpha).$
Moreover, in light of (i) of Corollary 1, we have $\sup_{F\in\mathcal{M}_{p}(m,v)}q_{\alpha}^{-}(F)=q_{\alpha}^{-}(F_{\mathcal{M}_{p}(m,v)}^{-})$. Hence, $q_{\alpha}^{-}(F_{\mathcal{M}_{p}(m,v)}^{-})=l(\alpha)$ for $\alpha\in(0,1)$. Note that $F_{\mathcal{M}_{p}(m,v)}^{-}(\cdot)$ is right-continuous. Hence, $F_{\mathcal{M}_{p}(m,v)}^{-}(x)=l_{+}^{-1}(x)$ for $x\in\mathbb{R}$. One can easily check that $l$ is strictly increasing and continuous over $(0,1)$. Thus we have that $F_{\mathcal{M}_{p}(m,v)}^{-}$ is continuous over $\mathbb{R}$, which further implies $\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}$ is continuous over $\mathbb{R}$.
Again, by Corollary 1 of Pesenti et al. (2020), we have
$\inf_{F\in\mathcal{M}_{p}(m,v)}q^{+}_{\alpha}(F)=m-v(1-\alpha)\left(\alpha^{p}(1-\alpha)+(1-\alpha)^{p}\alpha\right)^{-1/p}:=u(\alpha).$
Moreover, it follows from (i) of Corollary 1, we have $\inf_{F\in\mathcal{M}_{p}(m,v)}q_{\alpha}^{+}(F)=q_{\alpha}^{+}(F_{\mathcal{M}_{p}(m,v)}^{+})$. Combining the above two equations, we have $q_{\alpha}^{+}(F_{\mathcal{M}_{p}(m,v)}^{+})=u(\alpha)$ for $\alpha\in(0,1)$. Note that $u(\alpha)$ is strictly increasing and continuous over $(0,1)$. Hence, $F_{\mathcal{M}_{p}(m,v)}^{+}(x)=u_{+}^{-1}(x)$ for $x\in\mathbb{R}$ and $F_{\mathcal{M}_{p}(m,v)}^{+}$ is continuous.
For $\alpha\in(0,1)$ and $p=2$, we have $l(\alpha)=m+v\sqrt{{\alpha}/{(1-\alpha)}}$ and $u(\alpha)=m-v\sqrt{(1-\alpha)/\alpha}$. We obtain the desired results by computing $l_{+}^{-1}$ and $u_{+}^{-1}$.
Next, we show the attainability of $\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}$ and $F_{\mathcal{M}_{p}(m,v)}^{+}$. For any $x\in\mathbb{R}$, there exists $F_{k}\in\mathcal{M}_{p}(m,v)$ such that $\lim_{k\to\infty}F_{k}(x-)=\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}(x)$. By Chebyshev’s inequality, we have $\sup_{k\geqslant 1}(1-F_{k}(z))\leqslant\frac{v^{p}}{|z-\mu|^{p}}$ for $z>\mu$ and $\sup_{k\geqslant 1}F_{k}(z)\leqslant\frac{v^{p}}{|z-\mu|^{p}}$ for $z<\mu$. Hence, $\{F_{k},~{}k\geqslant 1\}$ is tight. It follows from Theorem 25.10 of Billingsley (1995)) that there exist a subsequence $\{F_{k_{l}},~{}l\geqslant 1\}$ and a distribution $F$ such that $F_{k_{l}}\to F$ weakly as $l\to\infty$. This implies $F_{k_{l}}^{-1}(t)\to F^{-1}(t)$ over $(0,1)$ a.e..
Next, we show $F\in\mathcal{M}_{p}(m,v)$. By Fatou’s lemma, we have
$\int_{0}^{1}|F^{-1}(t)-m|^{p}\mathrm{d}t=\int_{0}^{1}\liminf_{l\to\infty}|F_{k_{l}}^{-1}(t)-m|^{p}\mathrm{d}t\leqslant\liminf_{l\to\infty}\int_{0}^{1}|F_{k_{l}}^{-1}(t)-m|^{p}\mathrm{d}t\leqslant v^{p}.$
Using the fact $\sup_{l\geqslant 1}\int_{0}^{1}|F_{k_{l}}^{-1}(t)|^{p}\mathrm{d}t\leqslant(v+|m|)^{p}$, we have for $y>0$
$\int_{\{t:|F_{k_{l}}^{-1}(t)|\geqslant y\}}|F_{k_{l}}^{-1}(t)|\mathrm{d}t\leqslant\int_{\{t:|F_{k_{l}}^{-1}(t)|\geqslant y\}}|F_{k_{l}}^{-1}(t)|^{p}y^{1-p}\mathrm{d}t\leqslant(v+|m|)^{p}y^{1-p},$
which implies $\lim_{y\to\infty}\sup_{l\geqslant 1}\int_{\{t:|F_{k_{l}}^{-1}(t)|\geqslant y\}}|F_{k_{l}}^{-1}(t)|\mathrm{d}t=0.$ It means that $\{F_{k_{l}}^{-1},l\geqslant 1\}$ is uniformly integrable. Hence, we have $\int_{0}^{1}F^{-1}(t)\mathrm{d}t=\lim_{l\to\infty}\int_{0}^{1}F_{k_{l}}^{-1}(t)\mathrm{d}t=m$. Consequently, $F\in\mathcal{M}_{p}(m,v)$.
Note that there exists a sequence $\varepsilon_{n}\downarrow 0$ as $n\to\infty$ such that $F$ is continuous at all $x-\varepsilon_{n}$. Using the weak convergence of $F_{k_{l}}$ to $F$, we have $F(x-\varepsilon_{n})=\lim_{l\to\infty}F_{k_{l}}(x-\varepsilon_{n})\leqslant\lim_{l\to\infty}F_{k_{l}}(x-)=\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}(x)$. Letting $n\to\infty$, we have $F(x-)\leqslant\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}(x)$, which implies $F(x-)=\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}(x)$. Therefore, we conclude that $\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}$ is attainable.
For any $x\in\mathbb{R}$, there exists a sequence $F_{k}\in\mathcal{M}_{p}(m,v)$ such that $\lim_{k\to\infty}F_{k}(x)=F_{\mathcal{M}_{p}(m,v)}^{+}(x)$. Similarly as in the proof of the attainability of $\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}$, there exist a subsequence $\{F_{k_{l}},~{}l\geqslant 1\}$ and $F\in\mathcal{M}_{p}(m,v)$ such that $F_{k_{l}}\to F$ weakly as $l\to\infty$. Moreover, we could find $\varepsilon_{n}\downarrow 0$ such that $F$ is continuous at all $x+\varepsilon_{n}$. Using the fact that $F_{k_{l}}\to F$ weakly as $l\to\infty$, we have $F(x+\varepsilon_{n})=\lim_{l\to\infty}F_{k_{l}}(x+\varepsilon_{n})\geqslant\lim_{l\to\infty}F_{k_{l}}(x)=F_{\mathcal{M}_{p}(m,v)}^{+}(x)$. Letting $n\to\infty$, we obtain $F(x)\geqslant F_{\mathcal{M}_{p}(m,v)}^{+}(x)$, which implies $F(x)=F_{\mathcal{M}_{p}(m,v)}^{+}(x)$. Hence, $F_{\mathcal{M}_{p}(m,v)}^{+}$ is attainable. ∎
Combing Theorem 1, Propositions 3 and 6, we immediately arrive at the following result.
Theorem 2.
Let $F_{\mathcal{M}_{p}(m,v)}^{-}$ and $F_{\mathcal{M}_{p}(m,v)}^{+}$ be given by (5) and $\Lambda:\mathbb{R}\to[0,1]$. Then we have
  1. (i)
    $\sup_{F\in\mathcal{M}_{p}(m,v)}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}_{p}(m,v)}^{-})$; $\inf_{F\in\mathcal{M}_{p}(m,v)}q_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(F_{\mathcal{M}_{p}(m,v)}^{+})$; $\inf_{F\in\mathcal{M}_{p}(m,v)}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(F_{\mathcal{M}_{p}(m,v)}^{+})$;
  2. (ii)
    If $\Lambda$ is decreasing, then (i) remains true by replacing $\tilde{q}_{\Lambda}^{-}$ by $q_{\Lambda}^{-}$, $q_{\Lambda}^{+}$ by $\tilde{q}_{\Lambda}^{+}$ and $q_{\Lambda}^{-}$ by $\tilde{q}_{\Lambda}^{-}$; Moreover, $\sup_{F\in\mathcal{M}_{p}(m,v)}\tilde{q}_{\Lambda}^{+}(F)=\sup_{F\in\mathcal{M}_{p}(m,v)}q_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}_{p}(m,v)}^{-})$.
For the case $p=2$, we obtain the explicit formulas for $F^{-}_{\mathcal{M}_{2}(m,v)}$ and $F^{+}_{\mathcal{M}_{2}(m,v)}$. For general case $p\neq 2$, the values of $F^{-}_{\mathcal{M}_{(p,m,v)}}$ and $F^{+}_{\mathcal{M}_{(p,m,v)}}$ can be computed numerically using (5). For the particular case that $\Lambda$ is decreasing, it is more convenient to compute the robust $\Lambda$-quantiles using the expressions in Proposition 5 since we do not need to compute the inverse functions $l$ and $u$ to obtain $F_{\mathcal{M}_{p}{(m,v)}}^{-}$ and $F_{\mathcal{M}_{p}{(m,v)}}^{+}$.
In the following example, we compare the results for normal, exponential and uniform distributions with the worst or best values of $q_{\Lambda}^{-},q_{\Lambda}^{+},\tilde{q}_{\Lambda}^{-}$ and $\tilde{q}_{\Lambda}^{+}$, respectively.
Example 4.
Let $p=2$, $m=v=1$, and $\Lambda_{(\alpha,\beta;z)}:x\mapsto\alpha\mathds{1}_{\{x<z\}}+\beta\mathds{1}_{\{x\geqslant z\}}$. Using Theorem 2 and Proposition 5 , the numerical results of robust $\Lambda$-quantiles are displayed in Tables 1 and 2 for two different $\Lambda$ functions.
Table 1: Robust $\Lambda$-quantiles for an increasing $\Lambda$ with $\mathcal{M}=\mathcal{M}_{2}(m,v)$
$\Lambda_{(0.8,0.95;2)}$ $q_{\Lambda}^{-}$ $q_{\Lambda}^{+}$ $\tilde{q}_{\Lambda}^{-}$ $\tilde{q}_{\Lambda}^{+}$
$\mathrm{N}(1,1)$ 1.84 1.84 2.64 2.64
$\mathrm{exp}(1)$ 1.61 1.61 2.98 2.98
$U[1-\sqrt{3},1+\sqrt{3}]$ 2.57 2.57 2.57 2.57
Best case 0.51 0.51 0.51 0.51
Worst case 5.36 5.36 5.36 5.36
Table 2: Robust $\Lambda$-quantiles for a decreasing $\Lambda$ with $\mathcal{M}=\mathcal{M}_{2}(m,v)$
$\Lambda_{(0.95,0.8;2)}$ $q_{\Lambda}^{-}$ $q_{\Lambda}^{+}$ $\tilde{q}_{\Lambda}^{-}$ $\tilde{q}_{\Lambda}^{+}$
$\mathrm{N}(1,1)$ 2.00 2.00 2.00 2.00
$\mathrm{exp}(1)$ 2.00 2.00 2.00 2.00
$U[1-\sqrt{3},1+\sqrt{3}]$ 2.05 2.05 2.05 2.05
Best case 0.90 0.90 0.90 0.90
Worst case 3.00 3.00 3.00 3.00
For an increasing $\Lambda$, it can be observed from Table 1 that $q_{\Lambda}^{-}$ is strictly smaller than $\tilde{q}_{\Lambda}^{-}$, and $q_{\Lambda}^{+}$ is strictly smaller than $\tilde{q}_{\Lambda}^{+}$. This verifies the conclusion in (i) of Proposition 2. Note that the bold numbers in Table 1 may not be the true best-case or worst-case values; instead, they represent the lower bounds of the worst-case values or the upper bounds of the best-case values, respectively. Moreover, we observe that $\tilde{q}_{\Lambda}^{-}$, $\tilde{q}_{\Lambda}^{+}$, and the worst-case $\Lambda$-quantiles in Table 1 are consistently larger than those in Table 2. This shows that $\Lambda$-quantiles with an increasing $\Lambda$ tend to penalize more on scenarios with significant capital losses.
Our results in Theorem 2 can be applied to the robust portfolio selection. Let $\mathbf{X}=(X_{1},...,X_{n})\in\mathcal{X}^{n}$ represent the loss or negative return for $n$ different assets. The set of all possible portfolio positions are denoted by $\Delta_{n}=\left\{\mathbf{w}=\left(w_{1},\ldots,w_{n}\right)\in[0,1]^{n}:\sum_{i=1}^{n}w_{i}=1\right\}$. Note that here short-selling is not allowed. For $\mathbf{w}\in\Delta_{n}$, the risk of the portfolio $\mathbf{w}^{\top}\mathbf{X}$ is evaluated by $\rho\left(F_{\mathbf{w}^{\top}\mathbf{X}}\right)$, where $\rho$ is some risk measure. Suppose we know the mean and the upper bound of the covariance matrix of $\mathbf{X}$. Then the uncertainty set of the portfolio is given by
$\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)=\left\{F_{\mathbf{w}^{\top}\mathbf{X}}\in\mathcal{M}_{2}:\mathbb{E}[\mathbf{X}]=\boldsymbol{\mu},\operatorname{Cov}(\mathbf{X})\preceq\Sigma\right\},$ (6)
where $\Sigma$ is a semidefinite symmetric matrix and for a semidefinite symmetric matrix $B$, $B\preceq\Sigma$ means $\Sigma-B$ is positive semidefinite. It follows from the result of Popescu (2007) that $\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)=\mathcal{M}_{2}(\mathbf{w}^{\top}\boldsymbol{\mu},\sqrt{\mathbf{w}^{\top}\Sigma\mathbf{w}})$. The optimal portfolio selection with known mean and upper bound of covariance matrix is formulated as below
$\min_{\mathbf{w}\in{\Delta_{n}}}\sup_{F\in\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)}\rho\left(F\right).$
We refer to Pflug and Pohl (2018) for an overview of the portfolio selection with model uncertainty. Note that return of the portfolio can also be considered as a constraint in the above optimization problem. For instance, we can impose the constraint $\sum_{i=1}^{n}w_{i}\mu_{i}\leqslant c$ for some $c<0$ on the above optimization problem, requiring that the expected return of the portfolio is larger than $-c$. Then the optimization problem becomes $\min_{\mathbf{w}\in{\Delta_{n}^{\prime}}}\sup_{F\in\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)}\rho\left(F\right)$ with $\Delta_{n}^{\prime}=\left\{\mathbf{w}\in\Delta_{n}:\sum_{i=1}^{n}w_{i}\mu_{i}\leqslant c\right\}$, which does not change the nature of the problem. To simplify the problem, we here consider the portfolio selection problem with $c\geqslant\max_{i=1}^{n}\mu_{i}$.
Proposition 7.
Let $\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)$ be given in (6) and $\Lambda:\mathbb{R}\to[0,1]$. For $\rho=\tilde{q}_{\Lambda}^{+}$ with decreasing $\Lambda$ or $\rho=\tilde{q}_{\Lambda}^{-}$, we have
$\min_{\mathbf{w}\in{\Delta_{n}}}\sup_{F\in\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)}\rho\left(F\right)=\min_{\mathbf{w}\in{\Delta_{n}}}\rho\left(F^{-}_{\mathcal{M}_{2}(\mathbf{w}^{\top}\boldsymbol{\mu},\sqrt{\mathbf{w}^{\top}\Sigma\mathbf{w}})}\right),$ (7)
where
$F^{-}_{\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)}(x)=\frac{(x-\mathbf{w}^{\top}\boldsymbol{\mu})^{2}}{(\mathbf{w}^{\top}\Sigma\mathbf{w}+(x-\mathbf{w}^{\top}\boldsymbol{\mu})^{2})}\mathds{1}_{\{x\geqslant\mathbf{w}^{\top}\boldsymbol{\mu}\}}.$
Moreover, the optimal portfolio positions are given by the minimizer of the right-hand side of (7).
Proof.
According to the general projection property in Popescu (2007), the two sets $\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)$ and $\mathcal{M}_{2}(\mathbf{w}^{\top}\boldsymbol{\mu},\sqrt{\mathbf{w}^{\top}\Sigma\mathbf{w}})$ are identical. By Theorem 2, we obtain
$\sup_{F\in\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)}\rho\left(F\right)=\rho\left(F^{-}_{\mathcal{M}_{2}(\mathbf{w}^{\top}\boldsymbol{\mu},\sqrt{\mathbf{w}^{\top}\Sigma\mathbf{w}})}\right),$
which further implies
$\min_{\mathbf{w}\in{\Delta_{n}}}\sup_{F\in\mathcal{M}_{2}(\mathbf{w},\boldsymbol{\mu},\Sigma)}\rho\left(F\right)=\min_{\mathbf{w}\in{\Delta_{n}}}\rho\left(F^{-}_{\mathcal{M}_{2}(\mathbf{w}^{\top}\boldsymbol{\mu},\sqrt{\mathbf{w}^{\top}\Sigma\mathbf{w}})}\right).$
We complete the proof. ∎
Example 5.
We consider the case of $n=2$. Let $p=2$ and $\Lambda_{(\alpha,\beta;z)}:x\mapsto\alpha\mathds{1}_{\{x<z\}}+\beta\mathds{1}_{\{x\geqslant z\}}$. Assume that $\boldsymbol{\mu}_{1}=(0.5,1)^{\top}$, $\boldsymbol{\mu}_{2}=(5,6)^{\top}$, and the covariance matrices are given by
$\Sigma_{1}=\left(\begin{array}[]{cc}1&0.5\\ 0.5&1\end{array}\right),~{}~{}\text{and}~{}~{}\Sigma_{2}=\left(\begin{array}[]{cc}1&-0.5\\ -0.5&1\end{array}\right).$ (8)
Here, we consider only the worst case of $\tilde{q}^{-}_{\Lambda}$, as $\sup_{F\in\mathcal{M}_{p}(m,v)}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}_{p}(m,v)}^{-})$ holds for both increasing and decreasing $\Lambda$.
Figures 2 and 2 demonstrate the worst-case values of the portfolio as $w_{1}$ varies, showing that the optimal portfolio positions under the $\Lambda$-quantile criterion can differ from those under the VaR criterion. Additionally, for an increasing $\Lambda$ in Figure 2, when the mean of the portfolio is relatively small, the DM uses the small probability level of $0.8$ to determine the portfolio positions; when the mean of the portfolio is relatively large, a larger probability level of $0.95$ is used. But if $\Lambda$ is a decreasing function in Figure 2, this situation is reversed. This phenomenon is due to the fact that an increasing $\Lambda$ may penalize large losses, while a decreasing $\Lambda$ suggests that the DM accepts a smaller probability for larger potential losses, indicating a relative preference for risk. Moreover, when two assets are positively correlated, the worst-case values are higher compared to the case with negatively correlated assets, as negative correlation results in hedging effect.
Figure 1: The worst-case values of $\tilde{q}^{-}_{\Lambda}$ and $q^{-}_{\lambda}$ in $w_{1}$ with $\Lambda(\alpha,\beta;z)=\Lambda(0.8,0.95;3)$ (blue solid line) and $\lambda=0.9$ (black dashed line)
Figure 2: The worst-case values of $\tilde{q}^{-}_{\Lambda}$ and $q^{-}_{\lambda}$ in $w_{1}$ with $\Lambda(\alpha,\beta;z)=\Lambda(0.95,0.8;6)$ (blue solid line) and $\lambda=0.9$ (black dashed line)
4.2 Probability constraints via Wasserstein distance
The Wasserstein metric is a popular notion used in mass transportation and distributionally robust optimization; see e.g., Esfahani and Kuhn (2018) and Blanchet and Murthy (2019). In the one-dimensional setting, the Wasserstein metric has an explicit formula. For $p\geqslant 1$ and $F,G\in\mathcal{M}_{p}$, the $p$-Wasserstein distance between $F$ and $G$ is defined as
$W_{p}(F,G)=\left(\int_{0}^{1}\left|F^{-1}(u)-G^{-1}(u)\right|^{p}\mathrm{~{}d}u\right)^{1/p}.$ (9)
For $\varepsilon\geqslant 0$, the uncertainty set of an $\varepsilon$-Wasserstein ball around a baseline distribution $G\in\mathcal{M}_{p}$ is given by
$\mathcal{M}_{p}({G},\varepsilon)=\left\{F\in\mathcal{M}_{p}:W_{p}(F,{G})\leqslant\varepsilon\right\},$ (10)
where the parameter $\varepsilon$ represents the magnitude of uncertainty. Let $x_{+}=\max(x,0)$. In the following result, we will give the expressions of $F_{\mathcal{M}_{p}(G,\varepsilon)}^{-}$ and $F_{\mathcal{M}_{p}(G,\varepsilon)}^{+}$.
Proposition 8.
For $\varepsilon>0,p\geqslant 1$ and $G\in\mathcal{M}_{p}$, we have
$F_{\mathcal{M}_{p}(G,\varepsilon)}^{-}(x)=l_{+}^{-1}(x),~{}F_{\mathcal{M}_{p}(G,\varepsilon)}^{+}(x)=u_{+}^{-1}(x),~{}x\in\mathbb{R},$ (11)
where for $\alpha\in(0,1)$, $l(\alpha)$ is the unique solution of the following equation
$\int_{\alpha}^{1}\left(l(\alpha)-q^{-}_{t}(G)\right)_{+}^{p}\mathrm{~{}d}t=\varepsilon^{p},$
and $u(\alpha)$ is the unique solution of the following equation
$\int_{0}^{\alpha}\left(q_{t}^{-}(G)-u(\alpha)\right)_{+}^{p}\mathrm{~{}d}t=\varepsilon^{p}.$
Moreover, both $\widehat{F}_{\mathcal{M}_{p}(G,\varepsilon)}^{-}$ and $F_{\mathcal{M}_{p}(G,\varepsilon)}^{+}$ are continuous and attainable.
Proof.
By (iii) of Proposition 4 of Liu et al. (2022), we have for $\alpha\in(0,1)$, $\sup_{F\in\mathcal{M}_{p}(G,\varepsilon)}q^{-}_{\alpha}(F)=l(\alpha)$ is the unique solution of the following equation
$\int_{\alpha}^{1}\left(l(\alpha)-q^{-}_{t}(G)\right)_{+}^{p}\mathrm{~{}d}t=\varepsilon^{p}.$
By (i) of Corollary 1, we have $\sup_{F\in\mathcal{M}_{p}(G,\varepsilon)}q_{\alpha}^{-}(F)=q_{\alpha}^{-}(F_{\mathcal{M}_{p}(G,\varepsilon)}^{-})=l(\alpha)$. For $10<\alpha_{1}<\alpha_{2}<$ 1, it follows that $\int_{\alpha_{2}}^{1}\left(l(\alpha_{1})-q^{-}_{t}(G)\right)_{+}^{p}\mathrm{~{}d}t<\varepsilon^{p}$. This implies $l(\alpha_{2})>l(\alpha_{1})$. Hence, $l$ is strictly increasing over $(0,1)$, which implies $F_{\mathcal{M}_{p}(G,\varepsilon)}^{-}(x)=l_{+}^{-1}(x)$ for $x\in\mathbb{R}$ and $F_{\mathcal{M}_{p}(G,\varepsilon)}^{-}$ is continuous.
Next, we show the second statement. Note that $\inf_{F\in\mathcal{M}_{p}(G,\varepsilon)}q_{\alpha}^{+}(F)=-\sup_{F\in\mathcal{M}_{p}(\overline{G},\varepsilon)}q_{1-\alpha}^{-}(F):=u(\alpha)$, where $\overline{G}^{-1}(t)=-G^{-1}(1-t),~{}t\in(0,1)$. Using the above conclusion, we have $-u(\alpha)$ is the unique solution of
$\int_{1-\alpha}^{1}\left(-u(\alpha)+q^{-}_{1-t}(G)\right)_{+}^{p}\mathrm{~{}d}t=\varepsilon^{p}.$
Hence, $u(\alpha)$ is the unique solution of
$\int_{0}^{\alpha}\left(q^{-}_{t}(G)-u(\alpha)\right)_{+}^{p}\mathrm{~{}d}t=\varepsilon^{p}.$
It follows from (i) of Corollary 1 that $q_{\alpha}^{+}(F_{\mathcal{M}_{p}(G,\varepsilon)}^{+})=\inf_{F\in\mathcal{M}_{p}(G,\varepsilon)}q_{\alpha}^{+}(F)=u(\alpha)$. Moreover, for $\alpha_{1}<\alpha_{2}$, it follows that $\int_{0}^{\alpha_{2}}\left(q^{-}_{t}(G)-u(\alpha_{1})\right)_{+}^{p}\mathrm{~{}d}t>\varepsilon^{p}$, which implies $u(\alpha_{2})>u(\alpha_{1})$. Hence, $u$ is strictly increasing on $(0,1)$. Consequently, we have $F_{\mathcal{M}_{p}(G,\varepsilon)}^{+}(x)=u_{+}^{-1}(x),~{}x\in\mathbb{R}$ and $F_{\mathcal{M}_{p}(G,\varepsilon)}^{+}$ is continuous.
Next, we show the attainability of $\widehat{F}_{\mathcal{M}_{p}(G,\varepsilon)}^{-}$ and $F_{\mathcal{M}_{p}(G,\varepsilon)}^{+}$. For any $x\in\mathbb{R}$, there exists a sequence $F_{k}\in\mathcal{M}_{p}(G,\varepsilon)$ such that $\lim_{k\to\infty}F_{k}(x-)=\widehat{F}_{\mathcal{M}_{p}(G,\varepsilon)}^{-}(x)$. It follows that
$\int_{0}^{1}|F_{k}^{-1}(t)|^{p}\mathrm{d}t\leqslant\left(W_{p}(\delta_{0},G)+W_{p}(F_{k},G)\right)^{p}\leqslant\left(W_{p}(\delta_{0},G)+\varepsilon\right)^{p},$
where $\delta_{x}$ is the distribution with probability mass 1 on $x$. By Chebyshev’s inequality, we have $\sup_{k\geqslant 1}(1-F_{k}(x))\leqslant\frac{\left(W_{p}(\delta_{0},G)+\varepsilon\right)^{p}}{x^{p}}$ for $x>0$ and $\sup_{k\geqslant 1}F_{k}(x)\leqslant\frac{\left(W_{p}(\delta_{0},G)+\varepsilon\right)^{p}}{|x|^{p}}$ for $x<0$. Hence, $\{F_{k},~{}k\geqslant 1\}$ is tight. By Theorem 25.10 of Billingsley (1995)) that there exist a subsequence $\{F_{k_{l}},~{}l\geqslant 1\}$ and a distribution $F$ such that $F_{k_{l}}\to F$ weakly as $l\to\infty$, which implies $F_{k_{l}}^{-1}(t)\to F^{-1}(t)$ over $(0,1)$ a.s.. It follows from Fatou’s lemma that
$\int_{0}^{1}|F^{-1}(t)-G^{-1}(t)|^{p}\mathrm{d}t=\int_{0}^{1}\liminf_{l\to\infty}|F_{k_{l}}^{-1}(t)-G^{-1}(t)|^{p}\mathrm{d}t\leqslant\liminf_{l\to\infty}\int_{0}^{1}|F_{k_{l}}^{-1}(t)-G^{-1}(t)|^{p}\mathrm{d}t\leqslant\varepsilon^{p}.$
Hence, $F\in\mathcal{M}_{p}(G,\varepsilon)$. Using the same argument as in the proof of Proposition 6, we obtain $F(x-)=\widehat{F}_{\mathcal{M}_{p}(G,\varepsilon)}^{-}(x)$. Hence, $\widehat{F}_{\mathcal{M}_{p}(G,\varepsilon)}^{-}$ is attainable. Using the similar argument, we can show that $F_{\mathcal{M}_{p}(G,\varepsilon)}^{+}$ is attainable. The details are omitted. ∎
Following Theorem 1, Proposition 3, and Proposition 8, we immediately obtain the following conclusion.
Theorem 3.
For $\varepsilon>0,p\geqslant 1$, $G\in\mathcal{M}_{p}$, and $\Lambda:\mathbb{R}\to[0,1]$, let $F_{\mathcal{M}_{p}(G,\varepsilon)}^{-}$ and $F_{\mathcal{M}_{p}(G,\varepsilon)}^{+}$ be given by (11). Then we have
  1. (i)
    $\sup_{F\in\mathcal{M}_{p}(G,\varepsilon)}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}_{p}(G,\varepsilon)}^{-})$; $\inf_{F\in\mathcal{M}_{p}(G,\varepsilon)}q_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(F_{\mathcal{M}_{p}(G,\varepsilon)}^{+})$; $\inf_{F\in\mathcal{M}_{p}(G,\varepsilon)}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(F_{\mathcal{M}_{p}(G,\varepsilon)}^{+})$;
  2. (ii)
    If $\Lambda$ is decreasing, then (i) remains true by replacing $\tilde{q}_{\Lambda}^{-}$ by $q_{\Lambda}^{-}$, $q_{\Lambda}^{+}$ by $\tilde{q}_{\Lambda}^{+}$ and $q_{\Lambda}^{-}$ by $\tilde{q}_{\Lambda}^{-}$; Moreover, $\sup_{F\in\mathcal{M}_{p}(G,\varepsilon)}\tilde{q}_{\Lambda}^{+}(F)=\sup_{F\in\mathcal{M}_{p}(G,\varepsilon)}q_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}_{p}(G,\varepsilon)}^{-})$.
In the following example, we consider the uncertainty set $\mathcal{M}_{p}(G,\varepsilon)$ with $G$ being normal, exponential and uniform distributions. Note that $F_{\mathcal{M}_{p}(m,v)}^{-}$ and $F_{\mathcal{M}_{p}(m,v)}^{+}$ can be computed numerically using (11). For decreasing $\Lambda$, it is more convenient to calculate the best-case and worst case of $\Lambda$quantiles using the expressions in Proposition 5 as we do not need to compute the inverse functions of $l$ and $u$.
Example 6.
Let $p=1$, $\varepsilon=0.1$ and $\Lambda_{(\alpha,\beta;z)}:x\mapsto\alpha\mathds{1}_{\{x<z\}}+\beta\mathds{1}_{\{x\geqslant z\}}$. Applying the results in Theorem 3 and Proposition 5, we obtain the robust $\Lambda$-quantiles numerically, which are displayed in Tables 3 and 4. Again, the bold numbers in Table 3 represent the upper bounds for the best values and the lower bounds for the worst values.
We observe that the four $\Lambda$-quantiles in this example are all the same. Also, similar to Example 4, worst-case values of $\tilde{q}_{\Lambda}^{-}$ with an increasing $\Lambda$ are larger than those with a decreasing $\Lambda$. This indicates that $\tilde{q}_{\Lambda}^{-}$ with increasing $\Lambda$ functions are more conservative under this uncertainty set.
Table 3: Robust $\Lambda$-quantiles for an increasing $\Lambda$ with $\mathcal{M}=\mathcal{M}_{p}(G,\varepsilon)$
[…]
$\mathrm{N}(1,1)$ 1.90 1.90 1.90 1.90
$\mathrm{exp}(1)$ 2.00 2.00 2.00 2.00
$U[1-\sqrt{3},1+\sqrt{3}]$ 1.89 1.89 1.89 1.89
[…]
$\mathrm{N}(1,1)$ 5.06 5.06 5.06 5.06
$\sim\mathrm{exp}(1)$ 6.00 6.00 6.00 6.00
$U[1-\sqrt{3},1+\sqrt{3}]$ 4.65 4.65 4.65 4.65
Table 4: Robust $\Lambda$-quantiles for a decreasing $\Lambda$ with $\mathcal{M}=\mathcal{M}_{p}(G,\varepsilon)$
[…]
$\mathrm{N}(1,1)$ 1.06 1.06 1.06 1.06
$\mathrm{exp}(1)$ 2.11 2.11 2.11 2.11
$U[1-\sqrt{3},1+\sqrt{3}]$ 0.65 0.65 0.65 0.65
[…]
$\mathrm{N}(1,1)$ 2.90 2.90 2.90 2.90
$\mathrm{exp}(1)$ 3.11 3.11 3.11 3.11
$U[1-\sqrt{3},1+\sqrt{3}]$ 2.88 2.88 2.88 2.88
We next focus on the portfolio selection problem. For $p\geqslant 1$ and $a\geqslant 1$, the $p$-Wasserstein metric on $\mathbb{R}^{n}$ between $F,G\in\mathcal{M}_{p}\left(\mathbb{R}^{n}\right)$ is defined as
$W_{a,p}^{n}(F,G)=\inf_{F_{\mathbf{X}}=F,F_{\mathbf{Y}}=G}\left(\mathbb{E}\left[\|\mathbf{X}-\mathbf{Y}\|_{a}^{p}\right]\right)^{1/p},$
where $F_{\mathbf{X}}$ is the cdf of $\mathbf{X}$, and $\|\cdot\|_{a}$ is the $L^{a}$ norm on $\mathbb{R}^{n}$. If $n=1$, then $W_{a,p}^{d}$ is $W_{p}$ in (9) where the infimum is attained by comonotonicity via the Fréchet-Hoeffing inequality. Define the Wasserstein uncertainty set for a benchmark distribution $G\in\mathcal{M}_{p}(\mathbb{R}^{n})$ as
$\mathcal{M}_{a,p}^{n}(G,\varepsilon)=\left\{F\in\mathcal{M}_{p}\left(\mathbb{R}^{n}\right):W_{a,p}^{n}(F,G)\leqslant\varepsilon\right\},~{}\varepsilon>0.$
The univariate uncertainty set for the cdf of $\mathbf{w}^{\top}\mathbf{X}$ is denoted by
$\mathcal{M}_{\mathbf{w},a,p}(G,\varepsilon)=\left\{F_{\mathbf{w}^{\top}\mathbf{X}}:F_{\mathbf{X}}\in\mathcal{M}_{a,p}^{n}(G,\varepsilon)\right\},\quad G\in\mathcal{M}_{p}(\mathbb{R}^{d}).$
We next solve the following robust portfolio selection problem
$\min_{\mathbf{w}\in{\Delta_{n}}}\sup_{F\in\mathcal{M}_{\mathbf{w},a,p}(G,\varepsilon)}\rho\left(F\right).$ (12)
Lemma 1 (Theorem 5 of Mao et al. (2024)).
For $\varepsilon\geqslant 0,a\geqslant 1$ and $p\geqslant 1$, random vector $\mathbf{X}$ with $F_{\mathbf{X}}\in\mathcal{M}_{p}\left(\mathbb{R}^{n}\right)$ and $\mathbf{w}\in\mathbb{R}^{n}$, we have
$\mathcal{M}_{\mathbf{w},a,p}\left(F_{\mathbf{X}},\varepsilon\right)=\mathcal{M}_{p}\left(F_{\mathbf{w}^{\top}\mathbf{X}},\varepsilon\|\mathbf{w}\|_{b}\right),$
where $b$ satisfies $11/a+1/b=$ 1.
Based on Lemma 1, the optimization problems (12) can be solved using Theorem 3.
Proposition 9.
Suppose that $\varepsilon>0$, $p\geqslant 1$, $a\geqslant 1$, a random vector $\mathbf{X}$ satisfying $F_{\mathbf{X}}\in\mathcal{M}_{p}\left(\mathbb{R}^{n}\right)$ and $\Lambda:\mathbb{R}\to[0,1]$. For $\rho=\tilde{q}_{\Lambda}^{+}$ with decreasing $\Lambda$ or $\rho=\tilde{q}_{\Lambda}^{-}$, we have
$\min_{\mathbf{w}\in{\Delta_{n}}}\sup_{F\in\mathcal{M}_{\mathbf{w},a,p}(G,\varepsilon)}\rho\left(F\right)=\min_{\mathbf{w}\in{\Delta_{n}}}\rho\left(F^{-}_{\mathcal{M}_{p}\left(F_{\mathbf{w}^{\top}\mathbf{X}},\varepsilon\|\mathbf{w}\|_{a/(a-1)}\right)}\right),$ (13)
where $F^{-}_{\mathcal{M}_{p}\left(F_{\mathbf{w}^{\top}\mathbf{X}},\varepsilon\|\mathbf{w}\|_{a/(a-1)}\right)}(x)=u_{+}^{-1}(x),~{}x\in\mathbb{R}$ with $u(\alpha)$ being the unique solution of the following equation
$\int_{\alpha}^{1}\left(u(\alpha)-F_{\mathbf{w}^{\top}\mathbf{X}}^{-1}(s)\right)_{+}^{p}\mathrm{~{}d}s=(\varepsilon\|\mathbf{w}\|_{a/(a-1)})^{p},\quad\alpha\in(0,1).$
Moreover, the optimal portfolio positions are given by the minimizer of the right-hand side of (13).
Proof.
Using Theorem 3 and Lemma 1, we obtain the results immediately. ∎
Example 7.
Let the benchmark distribution $F_{0}=\textbf{t}(\tau,\boldsymbol{\mu}_{i},\Sigma_{i})$ ($i=1,2$), and denote by $F_{\tau}=\textbf{t}(\tau,0,1)$. Then for $x\in\mathbb{R}$, we have $F^{-}_{\mathcal{M}_{p}\left(F_{\mathbf{w}^{\top}\mathbf{X}},\varepsilon\|\mathbf{w}\|_{a/(a-1)}\right)}(x)=u_{+}^{-1}(x)$ with $u(\alpha)$ being the unique solution of the following equation
$\int_{\alpha}^{1}\left(u(\alpha)-\mathbf{w}^{\top}\boldsymbol{\mu}_{i}-\sqrt{\mathbf{w}^{\top}\Sigma_{i}\mathbf{w}}F_{\tau}^{-1}(s)\right)^{p}_{+}\mathrm{d}s=(\varepsilon\|\mathbf{w}\|_{a/(a-1)})^{p}.$
We consider the case of $n=2$. Let $p=1$, $a=2$, $\varepsilon=0.1$, $\tau=3$ and $\Lambda_{(\alpha,\beta;z)}:x\mapsto\alpha\mathds{1}_{\{x<z\}}+\beta\mathds{1}_{\{x\geqslant z\}}$. We assume that $\boldsymbol{\mu}_{1}=(0.5,1)^{\top}$, $\boldsymbol{\mu}_{2}=(3,4)^{\top}$, and the covariance matrices are given by (8).
Figures 4 and 4 show the worst-case values of the portfolio varying in $w_{1}$. The optimal portfolio positions under the $\Lambda\mathrm{VaR}$ criterion can be different from those under the VaR criterion. In Figure 4, with an increasing $\Lambda$, when the mean of the benchmark distribution is relatively small, the DM uses the small probability level of $0.8$ to determine the portfolio positions; when the mean of the portfolio is relatively large, a larger probability level of $0.95$ is used. But in Figure 4, where $\Lambda$ is a decreasing function, this situation is reversed. Additionally, the optimal portfolio positions are higher when the covariance matrix of the benchmark distribution is positively correlated compared to the case with negatively correlated assets. All these phenomena can be explained similarly to Example 5.
Figure 3: The worst-case values of $\tilde{q}^{-}_{\Lambda}$ and $q^{-}_{\lambda}$ in $w_{1}$ with $\Lambda(\alpha,\beta;z)=\Lambda(0.8,0.95;6)$ (blue solid line) and $\lambda=0.9$ (black dashed line)
Figure 4: The worst-case values of $\tilde{q}^{-}_{\Lambda}$ and $q^{-}_{\lambda}$ in $w_{1}$ with $\Lambda(\alpha,\beta;z)=\Lambda(0.95,0.8;6)$ (blue solid line) and $\lambda=0.9$ (black dashed line)
4.3 Marginal constraints in risk aggregation
For a loss vector $(X_{1},\dots,X_{n})$, suppose that the marginal distribution $X_{i}\sim F_{i}$ is known but the dependence structure is completely unknown. This assumption is motivated from the context where data from different correlated products are separately collected and thus no dependence information is available; see Embrechts et al. (2013) and Embrechts et al. (2015). All the possible distributions of the total loss $X_{1}+\dots+X_{n}$ are characterized by the aggregation set as below
$\mathcal{D}_{n}(\mathbf{F})=\left\{\text{ cdf of }X_{1}+\dots+X_{n}:X_{i}\sim F_{i},i=1,\ldots,n\right\},$
where $\mathbf{F}=\left(F_{1},\ldots,F_{n}\right)\in\mathcal{M}^{n}$. The worst-case and best-case values of $\rho$ over this uncertainty set are given by
$\overline{\rho}(\mathbf{F})=\sup\left\{\rho(G):G\in\mathcal{D}_{n}(\mathbf{F})\right\},~{}~{}\text{and}~{}~{}\underline{\rho}(\mathbf{F})=\inf\left\{\rho(G):G\in\mathcal{D}_{n}(\mathbf{F})\right\}.$
The worst-case value of quantiles over the aggregation set has been studied extensively in the literature. It is well known that worst-case value of quantiles in risk aggregation generally does not admit any analytical formula. Typically, the results in the literature either offer explicit expressions for some special type of marginal distributions or provide some upper bounds for the general marginal distributions. We refer to Embrechts et al. (2006) for bounds, and Wang and Wang (2016), Jakobsons et al. (2016) and Blanchet et al. (2020) for the analytical expressions.
Let $\mathcal{M}_{D}$ (resp. $\mathcal{M}_{I}$) denote the set of all distributions with decreasing (resp. increasing) densities. The expressions of the extreme probabilities over uncertainty set $\mathcal{D}_{n}(\mathbf{F})$ are displayed in the following proposition.
Proposition 10.
For $\mathbf{F}\in\mathcal{M}_{D}^{n}\cup\mathcal{M}_{I}^{n}$, we have
$F_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)=1-H_{\mathbf{F}}(x),~{}F_{\mathcal{D}_{n}(\mathbf{F})}^{+}(x)=H_{\overline{\mathbf{F}}}(-x),~{}x\in\mathbb{R},$ (14)
where
$H_{\mathbf{F}}(x)=\inf_{\mathbf{r}\in\Theta_{n}(x)}\left\{\sum_{i=1}^{n}\frac{1}{x-r}\int^{x-r+r_{i}}_{r_{i}}(1-F_{i}(y))\mathrm{d}y\right\}$
with $\mathbf{r}=(r_{1},\dots,r_{n})$, $r=\sum_{i=1}^{n}r_{i}$ and $\Theta_{n}=\{(r_{1},\dots,r_{n})\in\mathbb{R}^{n}:\sum_{i=1}^{n}r_{i}<x\},$ and $\overline{\mathbf{F}}=(\overline{F}_{1},\dots,\overline{F}_{n})$ with $\overline{F}_{i}(x)=1-F_{i}(-x),~{}x\in\mathbb{R}$. Moreover, both $\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}$ and $F_{\mathcal{D}_{n}(\mathbf{F})}^{+}$ are continuous and attainable.
Proof.
In light of Theorems 2 and 4 of Blanchet et al. (2020), we have
$\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{+}(F)=H_{\mathbf{F}}^{-1}(1-\alpha),~{}\alpha\in(0,1).$ (15)
Note that $\mathbf{F}\in\mathcal{M}_{D}^{n}\cup\mathcal{M}_{I}^{n}$ implies $F_{1}^{-1},\dots,F_{n}^{-1}$ are continuous over $(0,1)$. Hence, by Lemma 4.5 of Bernard et al. (2014), we have
$\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{-}(F)=\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{+}(F),~{}\alpha\in(0,1).$ (16)
Moreover, in light of (i) of Corollary 1, we have $\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{-}(F)=q_{\alpha}^{-}(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})$ for $\alpha\in(0,1)$. Consequently, $q_{\alpha}^{-}(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})=H_{\mathbf{F}}^{-1}(1-\alpha),~{}\alpha\in(0,1).$ As it is stated in the proof of Theorem 4 of Blanchet et al. (2020), $H_{\mathbf{F}}(x)$ is continuous over $\mathbb{R}$ and strictly decreasing over $(-\infty,\sum_{i=1}^{n}q_{1}^{-1}(F_{i}))$ and is equal to 00 on $[\sum_{i=1}^{n}q_{1}^{-1}(F_{i}),\infty)$. Hence, we have $F_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)=1-H_{\mathbf{F}}(x),~{}x\in\mathbb{R}$.
By (i) of Corollary 1, we have $\inf_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{+}(F)=q_{\alpha}^{+}(F_{\mathcal{D}_{n}(\mathbf{F})}^{+})$ for $\alpha\in(0,1)$. Direct computation gives $\inf_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{+}(F)=-\sup_{F\in\mathcal{D}_{n}(\overline{\mathbf{F}})}q_{1-\alpha}^{-}(F)$, where $\overline{\mathbf{F}}=(\overline{F}_{1},\dots,\overline{F}_{n})$ with $\overline{F}(x)=\lim_{y\downarrow x}1-F(-x),~{}x\in\mathbb{R}$. Note that $\overline{\mathbf{F}}\in\mathcal{M}_{D}^{n}\cup\mathcal{M}_{I}^{n}$. Hence, using (15) and (16), we have
$q_{\alpha}^{+}(F_{\mathcal{D}_{n}(\mathbf{F})}^{+})=-\sup_{F\in\mathcal{D}_{n}(\overline{\mathbf{F}})}q_{1-\alpha}^{-}(F)=-\sup_{F\in\mathcal{D}_{n}(\overline{\mathbf{F}})}q_{1-\alpha}^{+}(F)=-H_{\overline{\mathbf{F}}}^{-1}(\alpha),~{}\alpha\in(0,1).$
Note that $H_{\overline{\mathbf{F}}}(x)$ is strictly decreasing and continuous over $(-\infty,\sum_{i=1}^{n}q_{1}^{-1}(\overline{F}_{i}))$ and is equal to 00 on $[\sum_{i=1}^{n}q_{1}^{-1}(\overline{F}_{i}),\infty)$. Hence, we have $F_{\mathcal{D}_{n}(\mathbf{F})}^{+}(x)=H_{\overline{\mathbf{F}}}(-x),~{}x\in\mathbb{R}$.
Next, we show the attainability of $\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}$ and $F_{\mathcal{D}_{n}(\mathbf{F})}^{+}$. For any $x\in\mathbb{R}$, there exists a sequence of $F_{k}\in\mathcal{D}_{n}(\mathbf{F}),k\geqslant 1$ such that $\lim_{k\to\infty}F_{k}(x-)=\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)$. For each $F_{k}$, there exists a copula $C_{k}$ such that the joint distribution of $(X_{1},\dots,X_{n})$ is $C_{k}(F_{1},\dots,F_{n})$ and $X_{1}+\dots+X_{n}\sim F_{k}$. It follows from Theorem 25.10 of Billingsley (1995) that we could find a subsequence $\{C_{k_{m}},m\geqslant 1\}$ and a copula $C$ such that $C_{k_{m}}\to C$ weakly as $m\to\infty$, which implies $F_{k_{m}}\to F$ weakly as $m\to\infty$, where $(X_{1},\dots,X_{n})\sim C(F_{1},\dots,F_{n})$ and $X_{1}+\dots+X_{n}\sim F$. Hence, we have $F\in\mathcal{D}_{n}(\mathbf{F})$. Note that there exists a sequence of $y_{l}\uparrow x$ such that $F$ is continuous at all $y_{l}$. Hence, we have $F(y_{l})=\lim_{m\to\infty}F_{k_{m}}(y_{l})\leqslant\liminf_{m\to\infty}F_{k_{m}}(x-)=\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)$. Letting $l\to\infty$, it follows that $F(x-)\leqslant\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)$, implying the attainability of $\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}$. We can similarly show the attainability of $\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{+}$. ∎
Note that the expressions of the extreme probabilities in Proposition 14 are essentially the dual bounds in Theorem 4.17 of Rüschendorf (2013). In light of Theorem 1, Propositions 3 and 10, we obtain the following result.
Theorem 4.
For $\mathbf{F}\in\mathcal{M}_{D}^{n}\cup\mathcal{M}_{I}^{n}$ and $\Lambda:\mathbb{R}\to[0,1]$, let $F_{\mathcal{D}_{n}(\mathbf{F})}^{-}$ and $F_{\mathcal{D}_{n}(\mathbf{F})}^{+}$ be given by (14). Then we have
  1. (i)
    $\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})$; $\inf_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(F_{\mathcal{D}_{n}(\mathbf{F})}^{+})$; $\inf_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(F_{\mathcal{D}_{n}(\mathbf{F})}^{+})$;
  2. (ii)
    If $\Lambda$ is decreasing, then (i) remains true by replacing $\tilde{q}_{\Lambda}^{-}$ by $q_{\Lambda}^{-}$, $q_{\Lambda}^{+}$ by $\tilde{q}_{\Lambda}^{+}$ and $q_{\Lambda}^{-}$ by $\tilde{q}_{\Lambda}^{-}$; Moreover, $\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}\tilde{q}_{\Lambda}^{+}(F)=\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})$.
In what follows, we consider the portfolio selection problem, i.e., to choose a optimal portfolio position $\boldsymbol{w}\in\Delta_{n}$ such that
$R_{\rho}(\boldsymbol{w})=\rho\left(\sum_{i=1}^{n}w_{i}X_{i}\right)$
is minimized, where $\rho$ is a risk measure and $\left(X_{1},\ldots,X_{n}\right)$ represents the negative returns of $n$ different assets. Here we suppose $X_{1},\dots,X_{n}$ have the identical distribution $F$ as we aim to check whether diversification can reduce the risk.
Recall that a doubly stochastic matrix is a square matrix with nonnegative entries and the sum of each column/row is equal to 1. Let $\mathcal{Q}_{n}$ denote the set of all $n\times n$ doubly stochastic matrices. For two portfolio positions $\boldsymbol{w},\boldsymbol{\gamma}\in\Delta_{n}$, we can say that $\boldsymbol{\gamma}$ is more diversified than $\boldsymbol{w}$, denoted by $\boldsymbol{\gamma}\prec\boldsymbol{w}$, if $\boldsymbol{\gamma}=\Lambda\boldsymbol{w}$ for some $\Lambda\in\mathcal{Q}_{n}$. Note that this binary relationship is also called the majorization order; see e.g., Marshall (2011). Here, we suppose the marginal distribution $X_{i}\sim F$ is known but the dependence structure is completely unknown. Hence, we consider the worst-case scenario. That is to find the optimal portfolio position $\boldsymbol{w}\in\Delta_{n}$ such that
$\overline{R}_{\rho}(\boldsymbol{w})=\sup\left\{\rho\left(\sum_{i=1}^{n}w_{i}Y_{i}\right):Y_{1},\ldots,Y_{n}\sim F\right\}.$
is minimized. Note that in the above setup, we only minimize the risk but do not consider the return of the portfolio. That is because the expected return of the portfolio is a constant and any constraint on the expected return makes no real sense.
Proposition 11.
Suppose $\rho=\tilde{q}_{\Lambda}^{+}$ with decreasing $\Lambda$ or $\rho=\tilde{q}_{\Lambda}^{-}$, and $\left(X_{1},\ldots,X_{n}\right)$ has identical marginal distributions $F\in\mathcal{M}_{D}\cup\mathcal{M}_{I}$. If $\boldsymbol{\gamma}\prec\boldsymbol{w}$ , then $\overline{R}_{\rho}(\boldsymbol{\gamma})\geqslant\overline{R}_{\rho}(\boldsymbol{w})$.
Proof.
Write $\boldsymbol{\gamma}=\left(\gamma_{1},\ldots,\gamma_{n}\right)$ and $\boldsymbol{w}=\left(w_{1},\ldots,w_{n}\right)$. Take $X\sim F$, and let $\mathbf{F}$ and $\mathbf{G}$ be the tuples of marginal distributions of $\left(\gamma_{1}X,\ldots,\gamma_{n}X\right)$ and $\left(w_{1}X,\ldots,w_{n}X\right)$, respectively. It follows that
$\overline{R}_{\rho}(\boldsymbol{\gamma})=\overline{\rho}(\mathbf{F})\quad\text{ and }\quad\overline{R}_{\rho}(\boldsymbol{w})=\overline{\rho}(\mathbf{G}).$
Using $\boldsymbol{\gamma}\prec\boldsymbol{w}$, there exists $\Lambda\in\mathcal{Q}_{n}$ such that $(F_{1}^{-1},\dots,F_{n}^{-1})=\Lambda(G_{1}^{-1},\dots,G_{n}^{-1})$, denoted by $\mathbf{F}=\Lambda\otimes\mathbf{G}$. By Theorem 3 of Chen et al. (2022), for $\alpha\in(0,1)$, we have
$\sup_{G\in\mathcal{D}_{n}(\mathbf{G})}q_{\alpha}^{-}(G)\leqslant\sup_{G\in\mathcal{D}_{n}(\Lambda\otimes\mathbf{G})}q_{\alpha}^{-}(G)=\sup_{G\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{-}(G)$
for $\alpha\in(0,1)$. Hence, it follows from (i) of Corollary 1 that $q_{\alpha}^{-}(F_{\mathcal{D}_{n}(\mathbf{G})}^{-})\leqslant q_{\alpha}^{-}(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})$ for $\alpha\in(0,1)$. Using the fact that both $F_{\mathcal{D}_{n}(\mathbf{F})}^{-}$ and $F_{\mathcal{D}_{n}(\mathbf{G})}^{-}$ are right-continuous for $\alpha\in(0,1)$, we have $F_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)\geqslant F_{\mathcal{D}_{n}(\mathbf{G})}^{-}(x)$ for all $x\in\mathbb{R}$. By the monotonicity of $\rho$, we have $\rho(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})\leqslant\rho(F_{\mathcal{D}_{n}(\mathbf{G})}^{-})$. Moreover, by Theorem 4, we have $\overline{R}_{\rho}(\boldsymbol{\gamma})=\rho(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})$ and $\overline{R}_{\rho}(\boldsymbol{w})=\rho(F_{\mathcal{D}_{n}(\mathbf{G})}^{-})$. Hence, we have $\overline{R}_{\rho}(\boldsymbol{w})\leqslant\overline{R}_{\rho}(\boldsymbol{\gamma}).$
Note that Proposition 11 shows that for the assets with same marginal distributions $F\in\mathcal{M}_{D}\cup\mathcal{M}_{I}$, more diversified portfolio has higher risk under dependence uncertainty if $\rho=\tilde{q}_{\Lambda}^{+}$ with decreasing $\Lambda$ or $\rho=\tilde{q}_{\Lambda}^{-}$ is applied. This is not surprising as some similar conclusion is shown in Proposition 8 of Chen et al. (2022) for quantiles. Note that if $\Lambda$ is a constant, $\tilde{q}_{\Lambda}^{-}$ boils down to quantiles. Therefore, the conclusion in Proposition 11 can be viewed as an extension of Proposition 8 of Chen et al. (2022). Moreover, Proposition 11 also implies that under those assumptions, the optimal portfolio positions are the ones with only single asset.
5 Conclusion
This paper summaries some properties of $\Lambda$-quantiles and explores distributionally robust models of $\Lambda$-quantiles. We obtain robust $\Lambda$-quantiles on general uncertainty sets, showing that obtaining the robust $\Lambda$-quantiles relies on finding the extreme probabilities over the same uncertainty sets. This finding is surprising and significantly simplifies the problem, enabling us to utilize many existing results on the robust quantiles from the literature. We provide closed-form solutions for the uncertainty sets characterized by three different constraints: (i) moment constraints; (ii) probability constraints via Wasserstein distance; and (iii) marginal constraints in risk aggregation. Those results are applied to optimal portfolio selection under model uncertainty, and can also be extended to other problems such as optimal reinsurance, which will be discussed in future.
Acknowledgments
The research of Xia Han is supported by the National Natural Science Foundation of China (Grant No. 12301604, 12371471) and the Fundamental Research Funds for the Central Universities, Nankai University (Grant No. 63231138).